### v1

# Quantum effects and temperature dependence of the free energy barrier and shear frequency in bilayer graphene

Abstract
Little is known about the quantum effects of friction, and first-principles studies are virtually non-existent. In multilayer graphene, the order of the stacking of the layers plays a crucial role in the electronic properties and the manifestation of superconductivity, and friction affects how layers slide to change the stacking. Here, we study the effects of friction in bilayer graphene by focusing on quantum fluctuations in the free energy barrier that layers have to overcome to slide with respect to each other. We evaluate the free energy at the barrier in bilayer graphene through the self-consistent harmonic approximation, accounting for quantum and thermal ionic fluctuations. We obtain a very large reduction of the barrier of more than 30%. Temperature effects are weaker, although high temperatures induce a significant reduction of the barrier as well. Our approach is general and paves the way for systematically accounting for nuclear quantum and thermal effects in free energy barriers of other macroscopic systems.
There has been intense research on graphene since its discovery in 2004, and also on multilayer graphene systems. Among many other interesting properties, flat bands are present in multilayer rhombohedral (ABC-stacked) graphene (RG)Koshino (2010), and in fact, superconductivity has been observed in RG in 2021Zhou et al. (2021). Several works recently studied, both experimentally and theoretically, under which conditions RG can form Yankowitz et al. (2014); Geisenhof et al. (2019); Bouhafs et al. (2020); Gao et al. (2020); Li et al. (2020); Kerelsky et al. (2021). In particular, some of us unveiled how the stacking of graphene layers changes under shear stress Nery et al. (2020), and obtained that it produces many consecutive layers of RG. In fact, all samples with long-range RG that have been observed so far have been subject to some form of shear stress Lin et al. (2012); Henni et al. (2016); Henck et al. (2018); Yang et al. (2019); Shi et al. (2020); Bouhafs et al. (2020), and thus of friction.
To move graphene layers with respect to each other, they have to overcome a free energy barrier that separates different minima. When doing so, they experiment “stick-slip” motion: they move gradually with increasing shear stress, followed by sudden rearrangements and decrease of the stressLiu et al. (2012). The maximum shear-stress to “unlock” layers is correlated with the height of the barrier. Such barrier corresponds to the activation energy of a metastable state, where a graphene flake (from nanometers up to possibly micrometer scales) sitting on top of a large graphene layer is displaced relative to the equilibrium position. In clean samples with micrometer sized flakes however, different regions can have different stackings (AB and AC)Henck et al. (2018); Jiang et al. (2018), and a domain wall in the metastable state separates both regions. The displacement of layers in this case is more complex and is not considered in our work.
Superlubricity is another phenomenon present in systems involving graphene layers or graphite that has attracted a lot of attention(Liu et al., 2014; Sinclair et al., 2018), in which the energy landscape corresponding to the relative displacement of graphene layers is also central. Some works have studied how quantum effects, including zero-point renormalization (ZPR) and tunneling, affect the free-energy barrier of dissociative adsorptionMills and Jónsson (1994) and diffusionPoma et al. (2012); McIntosh et al. (2013); Cendagorta et al. (2016) of hydrogen. However, there are no studies of the quantum effects and temperature dependence of free-energy barriers in macroscopic systems.
A quantity closely related to the free energy barrier is the shear frequency, corresponding to the mode in which layers move rigidly in-plane and out of phase around the equilibrium position. The free energy barrier separates two equivalent minima, so the barrier and shear frequency are expected to have a similar temperature dependence (a lower curvature at the minima corresponds to a lower shear frequency and a lower free energy, at least close to the minima). The value of the shear frequency in graphite is a relatively well-established quantity, with experimental valuesLebedeva et al. (2017) oscillating between 42 - 45 cm-1, using for example inelastic X-ray scatteringMohr et al. (2007), neutron-coherent inelastic scatteringNicklow et al. (1972), Raman spectraTan et al. (2012); Hanfland et al. (1989), and coherent phonon spectroscopy Boschetto et al. (2013). In bilayer graphene, there are some experiments with values around 30 cm-1; for example, 28 $\pm$ 3 cm-1 in Refs. S Amelinckx, 1965; Boschetto et al., 2013, and 32 cm-1 in Ref. Tan et al., 2012. The temperature dependence of the shear modes, however, has been barely studied. Few experiments measure the temperature dependence of the shear mode, using Raman scattering in folded 2+2 (4 layer) grapheneCong and Yu (2014), or femtosecond pump-probe spectroscopy in graphiteMishina et al. (2000).
Another aspect of multilayer graphene and graphite that still needs further investigation and that serves as a consistency check on the adequacy of our approach is thermal expansion (TE). Ref. McQuade et al., 2021 measured it in bilayer and trilayer graphene for the first time. Ref. Mittal et al., 2021 uses QHA, which underestimates the out-of-plane TE at low temperaturesBailey and Yates (1970) and has a very rapid increase at high temperatures; in-plane, it also underestimates the TE at low temperaturesBailey and Yates (1970) and overestimates it at high temperaturesMorgan (1972). Much better out-of-plane results were obtained by using path integral molecular dynamics (PIMD)Herrero and Ramírez (2020), but while the TE flattens at 500 K in graphite, experimental values continue increasingMorgan (1972); Marsden et al. (2018). However, in-plane PIMD results of Ref. Herrero and Ramírez, 2020 vastly overestimate the TE.
Motivated by the aforementioned phenomena, especially by the lack of understanding of quantum effects in friction, we focus on studying the ZPR and temperature effects of the free-energy barrier. This is a very challenging problem due to the difficulty of determining the free energy in a non-equilibrium position. Here we tackle it using the state-of-the-art stochastic self-consistent harmonic approximation (SSCHA)Monacelli et al. (2021). First, we compare our results to the height of the barrier with fixed atoms. Then, using also the quasiharmonic approximation (QHA), we calculate the temperature dependence of the shear frequency and compare to the limited experimental data available. Finally, we determine the in-plane and out-of-plane TE coefficient in bilayer graphene and graphite, and compare our results to previous theoretical and experimental work.
We first focus on studying bilayer graphene, where the in-plane and (most importantly) out-of-plane interactions are given by the interatomic potential of Ref. Libbi et al., 2020 and Ref. Wen et al., 2018, respectively. The out-of-plane interactions determine the barrier. The energy profile of a layer moving with respect to the other one along the bond direction is given by Fig. 1, which gives good results when comparing to theory and experiment (see Table 1).
Table 1: Comparison of the barrier $V_{\mathrm{SP}}$ and the shear frequency at fixed ionic positions (classical) between the interatomic potential we used in this workLibbi et al. (2020); Wen et al. (2018), LDA, and theoretical calculations and experiments from previous works. The values are a little bit lower but close to that of adiabatic-connection fluctuation-dissipation theorem within the random phase approximation (ACFDT-RPA)Zhou et al. (2015) and experimental dataLebedeva et al. (2017).
2 layers
This work
2 layers
LDA
Previous
works
Small barrier $V_{\textrm{SP}}$
(meV/atom)
1.24 1.58 1.53 (RPA) Zhou et al. (2015)
Shear frequency
(cm-1) (bulk)
37 42 42-45 Lebedeva et al. (2017)
The lower layer is fixed in position A, and there are two minima of the same energy, AB and AC, at a distance of a bond length $d=1.42$. For layers to slide with respect to each other, they have to overcome a barrier $V_{\mathrm{SP}}$. SP is a saddle point in the full energy landscape, so the shear mode in the perpendicular direction (along the direction of two nearby AA maxima) is stable. In order to determine the ZPR and temperature dependence of the barrier, we use SSCHA to determine the free energy at the equilibrium position and SP.
The SSCHA minimizes the quantum free energy similarly to a variational Monte Carlo for nuclei, optimizing an ansatz density matrix of the most general Gaussian. The average atomic positions of the layers can be constrained at SP, and thanks to the simple shape of the variational ansatz, a quantum-free energy can be rigorously definedBianco et al. (2017); Monacelli et al. (2021). Notably, QHA cannot determine the barrier as the harmonic frequencies are not well-defined at SP. On the other hand, the SSCHA free energy landscape remains properly defined even in out-of-equilibrium and unstable configurations like SP. The Gaussian ansatz of the SSCHA allows separating the total free energy in an auxiliary harmonic contribution, which is long-ranged and can be interpolated up to a very fine $k$-mesh (40$\times$40), fundamental to avoid oversampling the energy of the shear mode, and a short-range anharmonic contribution, which can be properly converged already with small simulation cells. More details about QHA, SSCHA, and the interpolation method can be found in the Appendix.
Figure 1: Energy profile of bilayer graphene when moving one layer relative to the other along the bond direction. The lower layer is fixed in position A. Configuration AA (one layer on top of each other) is the least favorable configuration. There are two minima of the same energy, AB and AC, which correspond to the two stable configurations. To go from one minima to the other one applying shear stress, the upper layer has to overcome the barrier $V_{\mathrm{SP}}$, which has a value of 1.24 meV/atom relative to the minima. The diagrams below illustrate the different configurations, with the upper layer (darker blue) moving relative to the lower one in position A (lighter blue).
Figure 2: Barrier as a function of temperature. It is reduced from 1.24 meV/atom, the value at fixed nuclei (VSP in Fig. 1), to about 0.8 meV/atom, corresponding to a large ZPR of about 35%. The barrier does not change much up to about 500 K, but the value at 1000 K is about 20% smaller relative to the value at 100 K. The lines are a guide to the eye, as in the other figures of this work.
The barrier as a function of temperature can be seen in Fig. 2. The effects of the ZPR are very large: the barrier is reduced by an outstanding 35% with respect to the value at fixed nuclei. Since the frictional force is expected to be proportional to the barrier, quantum effects play a decisive role in the ability of layers to slide with respect to each other. Several other systems also present a reduction (sometimes also substantial) of the free energy relative to classical values, like adsorption or diffusion of hydrogen on several materialsMills and Jónsson (1994); Poma et al. (2012); McIntosh et al. (2013); Cendagorta et al. (2016). In terms of temperature effects, they are almost entirely covered by quantum fluctuations, since the barrier remains constant at around 0.8 meV/atom up to 500 K. Still when temperature increases up to 1000 K, the barrier is further reduced by 20% relative to the value at 100 K. Thus, high temperatures should help to reduce (static) friction between layers.
At the equilibrium position, one of the vibrational modes at $\Gamma$ is the shear mode in both directions of the plane (doubly degenerate, see Fig. 3). Its temperature dependence can be determined using both SSCHA and QHA. In the QHA, the TE is obtained by determining the lattice parameters that minimize the free energy at each temperature. The harmonic phonon frequencies calculated at such lattice parameters then also give temperature dependent frequencies. See Appendix Sec. A1.1 for more details. The third out of phase mode of the layers is the breathing mode, in the out-of-plane direction. A discussion of its temperature dependence can also be found in the Appendix.
The temperature dependence of the shear frequency can be seen in Fig. 4. The quantum effects evaluated within the SSCHA harden the shear frequency energy at low temperatures compared with the classical value at 0 K (dashed line). On the other hand, the ZPR in QHA softens the shear frequency. The temperature dependence of the shear mode is similar to the barrier, varying by about 20% from 0 to 1000 K, and having a larger dependence at temperatures above 500 K.
Figure 3: Bilayer graphene phonon dispersion. The inset shows a zoom in close to $\Gamma$ to better visualize the doubly degenerate shear modes S at $\Gamma$, and the breathing mode B. Due to the high speed velocity in graphene, the region of the shear mode in the BZ is small (a fraction of $\Gamma$M), and is overrepresented in small supercell calculations. The breathing mode, on the other hand, is flat, so nearby sampling points are not necessary to converge the free energy.
Regarding the experimental data, in Ref. Cong and Yu, 2014, the measurements are done in 2+2 folded graphene (4 layers). In order to compare with our 2-layer calculations, we multiply by a correction factor using a nearest-neighbor model for layersHe et al. (2016). In the case of bulk, such a factor is $\sqrt{2}$ (the force exerted by a layer below and above, proportional to the square of the frequency, is doubled when considering only nearest neighbor interactions), which we use to compare to the bulk measurements of Ref. Mishina et al., 2000. The temperature dependence of both SSCHA and QHA curves is similar to the experimental adjusted bulk values. Folded graphene values have a steeper dependence but are likely unsuitable for comparison, because layers fold with different rotational angles, while layers are aligned in standard graphite. The shear frequency has also been measured in other materials like bilayer NbSe2 in Ref. He et al., 2016, or bulk $h$-BN in Ref. Cuscó et al., 2019, which show a similar temperature dependence (adjusting the bulk value by a $\sqrt{2}$ factor) of roughly -0.5 cm-1/100 K.
Figure 4: Shear frequency as a function of temperature using SSCHA (blue circles) and QHA (red crosses). There is no experimental data for the temperature dependence of the shear frequency as a function of temperature for bilayer graphene. Experimental values correspond to bulk (blue crosses) and folded 2+2 graphene (orange circles), which are corrected by the factor of a nearest neighbor modelHe et al. (2016) to compare to the bilayer case (and rigidly shifted by -5 cm-1 to facilitate the visual comparison). The agreement with the temperature dependence of the bulk values is good.
To further check the adequacy of our method (the use of SSCHA with the in-planeLibbi et al. (2020) and out-of-planeWen et al. (2018) interatomic potentials), we look at the TE in the previous calculations and also in bulk graphite, and compare to experiments.
In Fig. 5, we show the TE coefficient calculated with SSCHA and QHA in bulk, compared with experimental dataBailey and Yates (1970); Marsden et al. (2018) and PIMD calculations Herrero and Ramírez (2020). SSCHA results underestimate the thermal expansion. PIMD results are overall very good, although they also give values somewhat lower than experimentsBailey and Yates (1970); Marsden et al. (2018).
In-plane SSCHA calculations give excellent results. QHA works well at low temperatures but departs from experiments at higher temperatures. PIMD calculations of Ref. Herrero and Ramírez, 2020, although they work very well out-of-plane, greatly overestimate in-plane TE, presumably due to the inadequacy of their interatomic potential. It would be interesting in future work to compare SSCHA and PIMD using the same potentials.
Figure 5: (a) Graphite out-of-plane coefficient of thermal expansion as a function of temperature using SSCHA and QHA. PIMD calculationsHerrero and Ramírez (2020) and experimental data Bailey and Yates (1970) are also included. PIMD overall agrees well, while QHA and SSCHA (with the potential of Ref. Wen et al., 2018) underestimates the TE. (b) Graphite in-plane coefficient of thermal expansion using SSCHA and QHA. QHA again underestimates the dependence at low temperatures, and also at high temperatures. On the other hand, PIMD gives a dependence that is surprisingly very far from experimental values. SSCHA gives excellent results both at low and high temperatures, considering also the disagreement within experimental data.
To conclude, we demonstrated that quantum fluctuations are fundamental to determining properties of layered materials that account for friction. Using the self-consistent harmonic approximation, which can calculate the quantum free energy at unstable configurations, and a novel interpolation method, we unveiled the suppression of the barrier separating stable configurations in bilayer graphene. Quantum fluctuations reduce the free energy barrier by 35%, and gets further reduced at high temperatures. The shear modes, which occupy a small region of phase space, have to be adequately sampled to correctly determine the free energy barrier as well as the thermal expansion. Although our approach underestimates the out-of-plane thermal expansion compared with experimental values, in-plane agreement with experiments is excellent. We also obtained that the temperature dependence of the barrier is similar to that of the shear frequency, which agrees well with existing measurements.
Acknowledgments
This work has received funding from the European Union’s Horizon 2020 research and innovation programme under Grant Agreement No. 881603, and from the MORE-TEM ERC-SYN project, Grant Agreement No. 951215.
Appendix
A1 Calculation details
A1.1 Quasi-harmonic approximation (QHA)
In the QHA, at each value of the lattice parameters $\{a_{i}\}$, the free energy $F(T,\{a_{i}\})=U(\{a_{i}\})-TS(T,\{a_{i}\})$ is given by the standard harmonic expression. That is,
$\begin{split}F(T,\{a_{i}\})=&U_{\mathrm{lat}}(\{a_{i}\})+\frac{1}{N}\sum_{\mathbf{q}\nu}\frac{1}{2}\hbar\omega_{\mathbf{q}\nu}(\{a_{i}\})+\\ &+\frac{1}{N}\sum_{\mathbf{q}\nu}kT\mathrm{ln}\left(1-e^{-\hbar\omega_{\mathbf{q}\nu}(\{a_{i}\})/kT}\right)\\ \end{split}$ (A1)
The value of the lattice parameters at a given temperature, $a_{i}(T)$, is determined by minimizing the free energy. Using an interatomic potential, the phonon frequencies at each value of the lattice parameters can be determined with little computational effort, and thus the QHA offers a quick method to determine the lattice expansion. Using first-principles calculations, phonons can be determined using density functional perturbation theory (DFPT), which uses a primitive cell and is much faster than frozen phonon (finite differences) calculfations in supercells. However, it does not account for the “true anharmonicity”Allen (2019), present at fixed values of the parameters when varying temperature. For example, it gives an increasing as opposed to decreasing temperature dependence of the G mode in graphene, due to the negative TE of the lattice parameter. Also, if the TE is negative and the lattice parameter reduces below the classical lattice value, as in the case of monolayer graphene at about 400 K (see later Fig. A10), then an acoustic mode becomes unstable and the QHA is ill-defined. These issues are not present in SSCHA.
A1.2 Stochastic self-consistent harmonic approximation (SSCHA)
In general, the free-energy of a ionic Hamiltonian $H=T+V$ is given by
$F_{H}=\mathrm{tr}(\rho_{H}H)+\frac{1}{\beta}\mathrm{tr}(\rho_{H}\mathrm{ln}\rho_{H})$ (A2)
where $\rho_{H}$ is the density matrix. In SSCHA, the density matrix is restricted to a trial harmonic Hamiltonian $\mathcal{H}=T+\mathcal{V}$, and the free energy is given by
$\mathcal{F}_{H}(\mathcal{H})=\mathrm{tr}(\rho_{\mathcal{H}}H)+\frac{1}{\beta}\mathrm{tr}(\rho_{\mathcal{H}}\mathrm{ln}\rho_{\mathcal{H}})$ (A3)
Due to a variational principle, $F_{H}\leq\mathcal{F}_{H}(\mathcal{H})$. The parameters of the harmonic Hamiltonian $\mathcal{H}$ include the centroids of the atoms, which correspond to the lattice parameters $a_{i}$ of the QHA, but it also includes the interatomic force constant matrix, which contains the information of all the phonon frequencies and accounts for the true anharmonicities. Thus, SSCHA has additional parameters, and is expected in general to give better results than QHA. Also, QHA does not satisfy a variational principle. It can be shown that
[…]
where $\mathbf{R}$ indicates a general ionic configuration and $\rho_{\mathcal{H}}(\mathbf{R})=\langle\mathbf{R}|e^{-\beta\mathcal{H}}|\mathbf{R}\rangle/Z_{\mathcal{H}}$ is the probability density of finding the crystal in a generic configuration $\mathbf{R}$ ($Z_{\mathcal{H}}=\mathrm{tr}[e^{-\beta\mathcal{H}}]$ the partition function). By stochastically sampling configurations $\mathbf{R}$ for a given density matrix $\rho_{\mathcal{H}}$, and determining the gradient of the free energy with respect the harmonic parameters (centroids and force constant matrix elements), the free energy can be minimized self-consistently. We used 24000 configurations in all calculations. A detailed description of SSCHA can be found for example in Refs. Errea et al., 2014 or Monacelli et al., 2021.
The logarithm term in Eq. A1 can be hard to converge for low frequencies. So rather than calculating $F_{\mathcal{H}}$ directly with the grid corresponding to the SC used in the minimization procedure, the force constant matrix can be Fourier interpolated in the usual way, and $F_{\mathcal{H}}$ can be calculated using a dense grid. This is the novel SSCHA interpolation method we mentioned in the main section, which was used with a $40\times 40$ in-plane grid to determine the barrier and also the out-of-plane TE. To properly determine the barrier, the interpolation scheme is used both at equilibrium and at SP, and then the interpolated free energies are subtracted.
The SSCHA frequencies are in principle auxiliary frequencies that minimize the free energy. In a purely harmonic system, such frequencies correspond to the physical frequencies. But in general, they have to be corrected with the so-called “bubble correction” (and also an additional higher order term in very anharmonic cases)Monacelli et al. (2021). Later in Fig. A5, we show the temperature dependendence of both the auxiliary and physical frequencies.
A2 Barrier
To determine the barrier, the free energy has to be determined at the equilibrium position and SP. At the equilibrium position, there are two shear modes, indicated with $\mathbf{S}$ in Fig. 3(b). As mentioned in the main text, the mode perpendicular to the bond direction shown in Fig. 1 is stable in SP, while the shear mode along the reaction coordinate (using the language of chemical reactions) is unstable. As opposed to the breathing mode, that is flat around $\Gamma$, the shear modes are only present in a small region of the BZ ($1/N^{2}$, with $N$ larger than 20), as a result of the large speed of sound in graphene. So in small supercells, the shear modes will be over-represented. This is more problematic at SP, where the SSCHA frequency corresponding to she unstable mode takes very small values (see Fig. A1), which makes the convergence of the free energy hard, mostly due to the logarithm term in $F_{\mathcal{H}}$ in Eq. (A4). Just determining the barrier with this equation and a $N\times N\times 1$ supercell, results in Fig. A2, which is not well converged.
In order to get better convergence, we use the interpolation method we mentioned in Sec. A1.2, and obtain a much more stable barrier as a function of $1/N$ (see Fig. A3). The $40\times 40$ interpolation grid used for $F_{\mathcal{H}}$ corresponds to a supercell of 6400 atoms. For 100, 200 and 300 K, the barrier seems well converged. For a 1000 K, the trend seems to indicate convergence within 0.02 meV/atom, so for the intermediate temperatures 400 and 500 K, the barrier is likely also converged within a similar error, resulting in Fig. 2 of the main text.
To better understand the reduction in the barrier, we show in Fig. A4 the SSCHA frequencies (around the shear modes) at equilibrium and SP at $T=100$ K . The SSCHA frequencies at equilibrium are similar to the standard harmonic values, as shown in the main text. The SSCHA shear frequencies at SP instead change significantly. The shear mode that looks like an acoustic mode takes a value of about 2 cm-1 at $\Gamma$ (see also Fig. A1). The rest of the SSCHA phonon dispersion presents small differences between equilibrium and SP frequencies.
Figure A1: Unstable shear frequency at SP as a function of $1/N$, for different temperatures. The fact that it does not stabilize shows that the shear mode is present in a small region of the BZ.
Figure A2: Barrier as a function of temperature, for different supercells, just using the $N\times N\times 1$ grid to determine the harmonic part of the free energy Eq. (A4). The barrier changes significantly with $N$, which means the shear mode is overrepresented. On the other hand, interpolating the harmonic part of the free energy, convergence is much better. See Fig. A3.
Figure A3: Barrier as a function of $1/N$, for different temperatures, interpolating the first term of Eq. A4. The barrier seems well converged at low temperatures and harder to converge at higher temperatures. Values are likely converged within less than 0.05 meV/atom.
Figure A4: SSCHA frequencies around the shear modes, at the equilibrium position (blue) and SP (red).
A3 Shear and breathing frequency
The SSCHA auxiliary and physical frequencies as a function of temperature can be observed in Fig. A5. The frequencies are not fully converged, and the converged value might be similar to the QHA. The bulk temperature dependence (corrected with the factor of the nearest neighbor model) agrees very well with QHA.
The other mode at $\Gamma$ that involves the layers moving out of phase is the breathing mode (out-of-plane instead of in-plane). Its temperature dependence can be seen in Fig. A6. It decreases more at low temperatures compared to the shear mode and barrier. There are no measurements of the breathing mode as a function of temperature, so our calculations serve as a prediction. The SSCHA curve is very well converged, so we expect it to have a temperature dependence that is closer to the experiment relative to the QHA curve. The auxiliary frequency for $N=8$ and 10 is almost the same (Fig. A7), so the physical values are also converged (within 0.5 cm-1 at 1000 K, and less for lower temperatures).
Figure A5: Shear frequency as a function of temperature, both for the physical frequency (full lines) and the auxiliary one (dashed). The experimental value is shifted and included twice, to facilitate the comparison of the temperature dependence with the QHA and SSCHA curves. The agreement of the experimental bulk value (adjusted with the nearest neighbor model to the bilayer case) with QHA is excellent, and with SSCHA is very good as well.
Figure A6: Breathing frequency as a function of temperature using SSCHA (black circles) and QHA (red crosses). The SSCHA calculations are well converged and serve as a prediction, since there is no experimental data in the literature.
Figure A7: Convergence of the auxiliary (dashed) and physical (full) breathing frequency. The physical frequency is well-converged at $N=10$.
A4 Thermal expansion (TE)
The interlayer distance and in-plane lattice parameter obtained in bulk with SSCHA can be seen in Fig. A8. The fit corresponds to Eq. (6) of Ref. Zorzi and Perottoni, 2021, both for in and out-of-plane TE ($b$ can be $a$ or $c_{d}$),
$b(T)=b_{0}\mathrm{exp}\left(\frac{X\theta}{e^{X/T}-1}\right),$ (A5)
which depends on parameters $b_{0}$, $X$ and $\theta$. In Fig. A9 we compare SSCHA and QHA for bilayer graphene and bulk. The TE for both SSCHA and QHA bulk is lower relative to bilayer. In-plane, Fig. A10, we can also see a clear reduction of the TE coefficient from monolayer to bulk. In what follows, we describe in detail convergence studies we did for SSCHA, and also for QHA calculations.
Figure A8: Out-of plane (a) and in-plane (b) SSCHA thermal expansion (blue dots). The fit corresponds to Eq. (A5).
Figure A9: SSCHA and QHA out-of-plane coefficient of TE for graphite and bilayer graphene.
Figure A10: In-plane coefficient of TE using SSCHA and QHA, for monolayer, bilayer and bulk.
A4.1 Bilayer
QHA. We first considered the out-of-plane TE by fixing $a$, and then we considered the full minimization of the free energy by varying both $a$ and $c_{d}$. In Fig. A11 we can see $c_{d}(T)$ both for fixed $a$, and for the full minimization. The full minimization has a slightly larger dependence with temperature, but fixing $a$ is a good approximation. Fig. A12 is analogous for $a(T)$, with a fixed value $c_{d}=3.5$. The curves are very similar, except for a small difference at high temperatures (of about 0.0005 Å at 1000 K).
We also studied convergence with the size of the grid (Fig. A13). At $N=12$, $c_{d}(T)$ is well converged (within 0.005 Å at 1000 K), although $N\sim 20$ to achieve full convergence. The shear mode increases thermal expansion because when layers are displaced relative to each other, the interlayer distance is larger (see Fig. A14).
In-plane, a larger $N$ of about 50 is needed to fully achieve convergence (A15).
[…]
On the other hand, $a(T)$ is already converged at about $N=8$ (Fig. A17).
A4.2 Bulk
QHA. In bulk, there is virtually no difference for the out-of-plane TE when fixing $a$, or considering also the minimization of $a$ (Fig. A13).
In-plane (Fig. A19), there is a small difference between both fixing or varying $a$, similar to the bilayer case.
SSCHA. Here we also used the interpolation method ($40\times 40$ grid), to go beyond the $N=8$ supercell that was possible in the direct SSCHA calculation, resulting in the blue dots of Fig. A8 a.
In-plane, based on $N=8$ already giving converged results in bilayer (Fig. A17), and also considering that the QHA convergence in bulk (Fig. A23) is much faster than in bilayer (Fig. A15), $N=8$ should give converged SSCHA results in bulk, which corresponds to the blue dots in Fig. A8 b.
Figure A11: Interlayer distance $c_{d}$ as a function of temperature in bilayer graphene using QHA, fixing the value of $a$ (dashed-blue) and doing the full minimization (full-black). In-plane TE has little effect on the out-of-plane TE.
Figure A12: Lattice parameter $a$ as a function of temperature in bilayer graphene using QHA, fixing the value of $c_{d}$ (dashed-blue) and doing the full minimization (full-black). In-plane TE has little effect on the out-of-plane TE.
Figure A13: Interlayer distance as a function of temperture in bilayer graphene, using QHA. At $N\sim 20$ $c_{d}$ is converged, although $N=12$ also gives a good result.
Figure A14: Interlayer distance profile along the bond direction, when moving a graphene layer relative to another one (analogous to Fig. 1, but for distance instead of energy). The minimum distance is at the equilibrium positions.
Figure A15: In-plane lattice parameter as a function of temperature in bilayer graphene for different $N$ using QHA. A large $N$ is needed to achieve convergence.
Figure A16: Interlayer distance as a function of temperature in bilayer graphene for different $N$ using SSCHA. Convergence is similar to the QHA case, Fig. A13, where $N=12$ gives a good value of $c_{d}(T)$.
Figure A17: In-plane lattice parameter as a function of temperature, for different $N$. At $N=8$ the curve is already well converged.
Figure A18: Interlayer distance as a function of temperature in bilayer graphene, using SSCHA, at two different values of the in-plane lattice parameter. The curves are similar to the ones of Fig. A11. In-plane TE has little effect on the out-of-plane TE.
Figure A19: Interlayer distance as a function of temperature in bulk using QHA, fixing the value of $c_{d}$ (dashed-blue) and doing the full minimization (full-black). In-plane TE has barely no effect on the out-of-plane TE.
Figure A20: In-plane lattice parameter as a function of temperature in bulk using QHA, fixing the value of $c_{d}$ (dashed-blue) and doing the full minimization (full-black). The out-of-plane TE has a slightly larger but similar effect relative to bilayer on the in-plane TE.
Figure A21: Interlayer distance as a function of temperature in graphite for different $N$ using QHA. Convergence is similar to the bilayer case.
Figure A22: Interlayer distance as a function of temperature in graphite for different $N_{z}$. $N_{z}=2$ is the value used in our calculations.
Figure A23: In-plane lattice parameter as a function of temperature in graphite using QHA, for different $N$. At $N=6$ convergence is already very good, as opposed to the bilayer case Fig. A16.

### v2

@@
 Here, we study the effects of friction in bilayer graphene by focusing on quantum fluctuations in the free energy barrier that layers have to overcome to slide with respect to each other.
-We evaluate the free energy at the barrier in bilayer graphene through the self-consistent harmonic approximation, accounting for quantum and thermal ionic fluctuations.
-We obtain a very large reduction of the barrier of more than 30%.
-Temperature effects are weaker, although high temperatures induce a significant reduction of the barrier as well.
+We evaluate the free energy at the barrier in bilayer graphene through the self-consistent harmonic approximation, accounting for quantum and thermal ionic fluctuations, and obtain a very large reduction of the barrier of more than 30% at low temperatures.
+There is little temperature dependence up to about 500 K, but it increases at higher temperatures.
 Our approach is general and paves the way for systematically accounting for nuclear quantum and thermal effects in free energy barriers of other macroscopic systems.
+I Introduction
 There has been intense research on graphene since its discovery in 2004, and also on multilayer graphene systems.
@@
 Superlubricity is another phenomenon present in systems involving graphene layers or graphite that has attracted a lot of attention(Liu et al., 2014; Sinclair et al., 2018), in which the energy landscape corresponding to the relative displacement of graphene layers is also central.
-Some works have studied how quantum effects, including zero-point renormalization (ZPR) and tunneling, affect the free-energy barrier of dissociative adsorptionMills and Jónsson (1994) and diffusionPoma et al.
+Yet another phenomena involving sliding surfaces is thermolubricity, which refers to the assistance from thermal excitations to overcome barriers.
+This term has been coined by some works that have studied the transition from stick-slip motion into stochastic fluctuations in friction at sufficiently high temperaturesBaykara et al.
+(2018).
+A few works have studied how quantum effects, including zero-point renormalization (ZPR) and tunneling, affect the free-energy barrier of dissociative adsorptionMills and Jónsson (1994) and diffusionPoma et al.
 (2012); McIntosh et al.
@@
 Ref.
-Mittal et al., 2021 uses QHA, which underestimates the out-of-plane TE at low temperaturesBailey and Yates (1970) and has a very rapid increase at high temperatures; in-plane, it also underestimates the TE at low temperaturesBailey and Yates (1970) and overestimates it at high temperaturesMorgan (1972).
+Mittal et al., 2021 uses the quasi-harmonic approxmimation (QHA), which underestimates the out-of-plane TE at low temperaturesBailey and Yates (1970) and has a very rapid increase at high temperatures; in-plane, it also underestimates the TE at low temperaturesBailey and Yates (1970) and overestimates it at high temperaturesMorgan (1972).
 Much better out-of-plane results were obtained by using path integral molecular dynamics (PIMD)Herrero and Ramírez (2020), but while the TE flattens at 500 K in graphite, experimental values continue increasingMorgan (1972); Marsden et al.
@@
 (2021).
-First, we compare our results to the height of the barrier with fixed atoms.
-Then, using also the quasiharmonic approximation (QHA), we calculate the temperature dependence of the shear frequency and compare to the limited experimental data available.
+First, in Sec.II, we briefly introduce SSCHA and the QHA, and explain how is the barrier determined.
+Then, in Sec.
+III, we present our main results and compare to the height of the barrier with fixed atoms.
+Subsequently, using also the QHA, we calculate the temperature dependence of the shear frequency and compare to the limited experimental data available.
 Finally, we determine the in-plane and out-of-plane TE coefficient in bilayer graphene and graphite, and compare our results to previous theoretical and experimental work.
+In Sec.
+IV, we summarize the conclusions of our work.
+II Theoretical and Computational Framework
+Here, we briefly describe the QHA, SSCHA, and a new SSCHA interpolation method used to determine the barrier.
+We also introduce the equation that is used to fit SSCHA thermal expansion results.
+A more detailed description of SSCHA can be found in Refs.
+Errea et al., 2014 or Monacelli et al., 2021.
+II.1 Quasi-harmonic approximation (QHA)
+In the QHA, at each value of the lattice parameters $\{a_{i}\}$, the free energy $F(T,\{a_{i}\})=U(\{a_{i}\})-TS(T,\{a_{i}\})$ is given by the standard harmonic expression. That is,
+$\begin{split}F(T,\{a_{i}\})=&U_{\mathrm{lat}}(\{a_{i}\})+\frac{1}{N}\sum_{\mathbf{q}\nu}\frac{1}{2}\hbar\omega_{\mathbf{q}\nu}(\{a_{i}\})+\\ &+\frac{1}{N}\sum_{\mathbf{q}\nu}kT\mathrm{ln}\left(1-e^{-\hbar\omega_{\mathbf{q}\nu}(\{a_{i}\})/kT}\right).\\ \end{split}$ (1)
+The value of the lattice parameters at a given temperature, $a_{i}(T)$, is determined by minimizing the free energy.
+Using an interatomic potential, the phonon frequencies at each value of the lattice parameters can be determined with little computational effort, and thus the QHA offers a quick method to determine the lattice expansion.
+Using first-principles calculations, phonons can be determined using density functional perturbation theory (DFPT), which uses a primitive cell and is much faster than frozen phonon (finite differences) calculations in supercells.
+However, it does not account for the “true anharmonicity”Allen (2019), present at fixed values of the parameters when varying temperature.
+For example, it gives an increasing as opposed to decreasing temperature dependence of the G mode in graphene, due to the negative TE of the lattice parameter.
+Also, if the TE is negative and the lattice parameter reduces below the classical lattice value, as in the case of monolayer graphene at about 400 K (see later Fig.
+A10), then an acoustic mode becomes unstable and the QHA is ill-defined.
+These issues are not present in SSCHA.
+II.2 Stochastic self-consistent harmonic approximation (SSCHA)
+In general, the free-energy of a ionic Hamiltonian $H=T+V$ is given by
+$F_{H}=\mathrm{tr}(\rho_{H}H)+\frac{1}{\beta}\mathrm{tr}(\rho_{H}\mathrm{ln}\rho_{H}),$ (2)
+where $\rho_{H}$ is the density matrix. In SSCHA, the density matrix is restricted to a trial harmonic Hamiltonian $\mathcal{H}=T+\mathcal{V}$, and the free energy is given by
+$\mathcal{F}_{H}(\mathcal{H})=\mathrm{tr}(\rho_{\mathcal{H}}H)+\frac{1}{\beta}\mathrm{tr}(\rho_{\mathcal{H}}\mathrm{ln}\rho_{\mathcal{H}}).$ (3)
+Due to a variational principle, $F_{H}\leq\mathcal{F}_{H}(\mathcal{H})$.
+The parameters of the harmonic Hamiltonian $\mathcal{H}$ include the centroids of the atoms, which correspond to the lattice parameters $a_{i}$ of the QHA, but it also includes the interatomic force constant matrix, which contains the information of all the phonon frequencies and accounts for the true anharmonicities.
+Thus, SSCHA has additional parameters, and is expected in general to give better results than QHA.
+Also, QHA does not satisfy a variational principle.
+It can be shown that
+[…]
+where $\mathbf{R}$ indicates a general ionic configuration and $\rho_{\mathcal{H}}(\mathbf{R})=\langle\mathbf{R}|e^{-\beta\mathcal{H}}|\mathbf{R}\rangle/Z_{\mathcal{H}}$ is the probability density of finding the crystal in a generic configuration $\mathbf{R}$ ($Z_{\mathcal{H}}=\mathrm{tr}[e^{-\beta\mathcal{H}}]$ is the partition function).
+By stochastically sampling configurations $\mathbf{R}$ for a given density matrix $\rho_{\mathcal{H}}$, and determining the gradient of the free energy with respect the harmonic parameters (centroids and force constant matrix elements), the free energy can be minimized self-consistently.
+Thanks to the variational ansatz, the SSCHA quantum free energy is rigorously defined, even in out-of-equilibrium and unstable configurations like SP.
+Notably, on the other hand, the QHA cannot determine the barrier, since the harmonic frequencies are not well defined at SP.
+The logarithm term in the harmonic term Eq.
+(1) can be hard to converge for low frequencies.
+Separating the total free energy as in Eq.
+(4) allows to interpolate the auxiliary harmonic term up to a very fine $q$-mesh, which is fundamental to avoid oversampling the energy of the long-ranged shear modes, while the anharmonic contribution can be properly converged already with small simulation cells.
+So rather than calculating $F_{\mathcal{H}}$ directly with the grid corresponding to the SC used in the minimization procedure, the force constant matrix can be Fourier interpolated in the usual way, and $F_{\mathcal{H}}$ can be calculated using a dense mesh.
+This procedure is followed both at SP and the equilibrium position, and the barrier is obtained by taking the difference of the corresponding free energies.
+More technical details are provided in the Appendix, Sec.
+A1.
+The SSCHA frequencies are in principle auxiliary frequencies that minimize the free energy.
+In a purely harmonic system, such frequencies correspond to the physical frequencies.
+But in general, they have to be corrected with the so-called “bubble correction” (and also an additional higher order term in very anharmonic cases)Monacelli et al.
+(2021), which we do in Sec.
+III.2 to obtain the temperature dependence of the shear frequency.
+Convergence with the supercell size of both the physical and auxiliary frequencies can be seen in the Appendix, Fig.
+A5.
+II.3 Thermal expansion (TE)
+As opposed to the barrier, the thermal expansion (TE) can also be calculated using the QHA, which although is much faster, is also in principle less accurate.
+With both QHA and SSCHA we determine the in-plane and out-of-plane lattice parameters as a function of temperature, $a(T)$ and $c(T)$, to then determine the in-plane and out-of-plane coefficients of TE, $\alpha_{\mathrm{a}}(T)=\tfrac{1}{a}\tfrac{da}{dT}$ and $\alpha_{\mathrm{c}}(T)=\tfrac{1}{c}\tfrac{dc}{dT}$.
+In the QHA case, the TE can calculated quickly and accurately for many temperatures, so we just use finite differences.
+For SSCHA, to reduce numerical error, we use Eq.
+(6) of Ref.
+Zorzi and Perottoni, 2021,
+$b(T)=b_{0}\mathrm{exp}\left(\frac{X\theta}{e^{X/T}-1}\right),$ (5)
+which depends on parameters $b_{0}$, $X$ and $\theta$, to fit $a(T)$ and $c_{d}(T)$ ($b$ can be $a$ or $c_{d}$). Then we determine the coefficients of TE from the derivative of the fit.
+III Results
+III.1 Barrier
 We first focus on studying bilayer graphene, where the in-plane and (most importantly) out-of-plane interactions are given by the interatomic potential of Ref.
@@
 The energy profile of a layer moving with respect to the other one along the bond direction is given by Fig.
-1, which gives good results when comparing to theory and experiment (see Table 1).
-Table 1: Comparison of the barrier $V_{\mathrm{SP}}$ and the shear frequency at fixed ionic positions (classical) between the interatomic potential we used in this workLibbi et al.
+1, which gives good reults when comparing to theory and experiment (see Table 1).
+Table 1: Comparison of the barrier $V_{\mathrm{SP}}$ and the shear frequency at fixed ionic positions (classical at $T=0$ K) between the interatomic potential we used in this workLibbi et al.
 (2020); Wen et al.
@@
 SP is a saddle point in the full energy landscape, so the shear mode in the perpendicular direction (along the direction of two nearby AA maxima) is stable.
-In order to determine the ZPR and temperature dependence of the barrier, we use SSCHA to determine the free energy at the equilibrium position and SP.
-The SSCHA minimizes the quantum free energy similarly to a variational Monte Carlo for nuclei, optimizing an ansatz density matrix of the most general Gaussian.
-The average atomic positions of the layers can be constrained at SP, and thanks to the simple shape of the variational ansatz, a quantum-free energy can be rigorously definedBianco et al.
-(2017); Monacelli et al.
-(2021).
-Notably, QHA cannot determine the barrier as the harmonic frequencies are not well-defined at SP.
-On the other hand, the SSCHA free energy landscape remains properly defined even in out-of-equilibrium and unstable configurations like SP.
-The Gaussian ansatz of the SSCHA allows separating the total free energy in an auxiliary harmonic contribution, which is long-ranged and can be interpolated up to a very fine $k$-mesh (40$\times$40), fundamental to avoid oversampling the energy of the shear mode, and a short-range anharmonic contribution, which can be properly converged already with small simulation cells.
-More details about QHA, SSCHA, and the interpolation method can be found in the Appendix.
+In order to determine the ZPR and temperature dependence of the barrier, we use SSCHA to determine the free energy at the equilibrium position and SP, using the interpolation method described in Sec.
+II.2.
 Figure 1: Energy profile of bilayer graphene when moving one layer relative to the other along the bond direction.
@@
 It is reduced from 1.24 meV/atom, the value at fixed nuclei (VSP in Fig.
-1), to about 0.8 meV/atom, corresponding to a large ZPR of about 35%.
+1), to about 0.8 meV/atom, corresponding to a large reduction of about 35% due to anharmonicities.
 The barrier does not change much up to about 500 K, but the value at 1000 K is about 20% smaller relative to the value at 100 K.
 The lines are a guide to the eye, as in the other figures of this work.
-The barrier as a function of temperature can be seen in Fig.
-2.
-The effects of the ZPR are very large: the barrier is reduced by an outstanding 35% with respect to the value at fixed nuclei.
-Since the frictional force is expected to be proportional to the barrier, quantum effects play a decisive role in the ability of layers to slide with respect to each other.
-Several other systems also present a reduction (sometimes also substantial) of the free energy relative to classical values, like adsorption or diffusion of hydrogen on several materialsMills and Jónsson (1994); Poma et al.
-(2012); McIntosh et al.
-(2013); Cendagorta et al.
-(2016).
-In terms of temperature effects, they are almost entirely covered by quantum fluctuations, since the barrier remains constant at around 0.8 meV/atom up to 500 K.
-Still when temperature increases up to 1000 K, the barrier is further reduced by 20% relative to the value at 100 K.
+[…]
+To confirm that quantum fluctuations are very minor down to 100 K, we performed SSCHA calculations in the classical limit $\hbar\omega/kT\rightarrow 0$ (which recovers the classical Boltzmann distribution for the distorted configurations) at 100 K and 300 K.
+We obtained comparable results, except for a small difference of less than 0.05 meV/atom at 100 K with respect to the quantum values.
+Much lower temperatures would have to be considered to recover the classical value at fixed ions.
+Thus, it is the anharmonic vibrations that play a decisive role in the ability of layers to slide with respect to each other.
+In a previous workNery et al.
+(2020), using density functional theory with an LDA functional (see Table 1), some of us obtained that the maximum shear stress in simulated stick-slip motion is 0.2 GPa, and we mentioned a good agreement with experimental values 0.14 GPa (Ref.
+Liu et al., 2012) and 0.1 GPa (Ref.
+Dienwiebel et al., 2004).
+Assuming the maximum shear stress decreases by 35% just like the barrier, we obtain 0.13 GPa, in excellent agreement with experiments.
+Other systems involving hydrogen diffusion across a free energy barrier do present significant differences between classical and quantum free energy barriers due to the large quantum fluctuations of hydrogen.
+Such differences have been observed at 100 K (Ref.
+Mills and Jónsson, 1994), 250 K (Ref.
+McIntosh et al., 2013), and even higher temperaturesPoma et al.
+(2012), although other worksCendagorta et al.
+(2016) obtain very similar barriers at already 200 K.
+Regarding the evolution of the barrier as the temperature increases, there is a weak dependence up to about 500 K, since the barrier remains close to around 0.8 meV/atom.
+However, when temperature increases up to 1000 K, the barrier is further reduced by 20% relative to the value at 100 K.
 Thus, high temperatures should help to reduce (static) friction between layers.
+In particular, it could aid the production of ABC-stacked multilayer graphene from its more common AB-stacked version through shear stress, as mentioned in the Introduction.
+More in general, high temperatures could also help engineer specific stacking arrangements in other layered materials.
+III.2 Shear frequency
 At the equilibrium position, one of the vibrational modes at $\Gamma$ is the shear mode in both directions of the plane (doubly degenerate, see Fig.
@@
 In the QHA, the TE is obtained by determining the lattice parameters that minimize the free energy at each temperature.
-The harmonic phonon frequencies calculated at such lattice parameters then also give temperature dependent frequencies.
-See Appendix Sec.
-A1.1 for more details.
-The third out of phase mode of the layers is the breathing mode, in the out-of-plane direction.
-A discussion of its temperature dependence can also be found in the Appendix.
+The third out of phase mode of the layers is the breathing mode, in the out-of-plane direction. A discussion of its temperature dependence can be found in the Appendix.
 The temperature dependence of the shear frequency can be seen in Fig.
@@
 On the other hand, the ZPR in QHA softens the shear frequency.
-The temperature dependence of the shear mode is similar to the barrier, varying by about 20% from 0 to 1000 K, and having a larger dependence at temperatures above 500 K.
+The temperature dependence of the shear mode is similar to that of the barrier, varying by about 20% from 0 to 1000 K, and having a larger dependence at temperatures above 500 K.
 Figure 3: Bilayer graphene phonon dispersion.
@@
 The agreement with the temperature dependence of the bulk values is good.
+III.3 Thermal expansion
 To further check the adequacy of our method (the use of SSCHA with the in-planeLibbi et al.
@@
 (2018) and PIMD calculations Herrero and Ramírez (2020).
+SSCHA curves result from fitting Eq.
+(5) to $a(T)$ and $c(T)$ (see Fig.
+A8) and then taking the derivative.
 SSCHA results underestimate the thermal expansion.
@@
 SSCHA gives excellent results both at low and high temperatures, considering also the disagreement within experimental data.
-To conclude, we demonstrated that quantum fluctuations are fundamental to determining properties of layered materials that account for friction.
+IV Conclusions
+To conclude, we demonstrated that anharmonicities are fundamental to determining properties of layered materials that account for friction.
 Using the self-consistent harmonic approximation, which can calculate the quantum free energy at unstable configurations, and a novel interpolation method, we unveiled the suppression of the barrier separating stable configurations in bilayer graphene.
-Quantum fluctuations reduce the free energy barrier by 35%, and gets further reduced at high temperatures.
+Anharmonicities reduce the free energy barrier by 35%, and gets further reduced at high temperatures.
+On the other hand, quantum fluctuations are not important down to 100 K.
 The shear modes, which occupy a small region of phase space, have to be adequately sampled to correctly determine the free energy barrier as well as the thermal expansion.
@@
 951215.
+J.P.N.
+is currently supported by the European Union under a Marie Skłodowska-Curie Individual Fellowship, Project GreenNP No.
+101151380.
 Appendix
-A1 Calculation details
-A1.1 Quasi-harmonic approximation (QHA)
-In the QHA, at each value of the lattice parameters $\{a_{i}\}$, the free energy $F(T,\{a_{i}\})=U(\{a_{i}\})-TS(T,\{a_{i}\})$ is given by the standard harmonic expression. That is,
-$\begin{split}F(T,\{a_{i}\})=&U_{\mathrm{lat}}(\{a_{i}\})+\frac{1}{N}\sum_{\mathbf{q}\nu}\frac{1}{2}\hbar\omega_{\mathbf{q}\nu}(\{a_{i}\})+\\ &+\frac{1}{N}\sum_{\mathbf{q}\nu}kT\mathrm{ln}\left(1-e^{-\hbar\omega_{\mathbf{q}\nu}(\{a_{i}\})/kT}\right)\\ \end{split}$ (A1)
-The value of the lattice parameters at a given temperature, $a_{i}(T)$, is determined by minimizing the free energy.
-Using an interatomic potential, the phonon frequencies at each value of the lattice parameters can be determined with little computational effort, and thus the QHA offers a quick method to determine the lattice expansion.
-Using first-principles calculations, phonons can be determined using density functional perturbation theory (DFPT), which uses a primitive cell and is much faster than frozen phonon (finite differences) calculfations in supercells.
-However, it does not account for the “true anharmonicity”Allen (2019), present at fixed values of the parameters when varying temperature.
-For example, it gives an increasing as opposed to decreasing temperature dependence of the G mode in graphene, due to the negative TE of the lattice parameter.
-Also, if the TE is negative and the lattice parameter reduces below the classical lattice value, as in the case of monolayer graphene at about 400 K (see later Fig.
-A10), then an acoustic mode becomes unstable and the QHA is ill-defined.
-These issues are not present in SSCHA.
-A1.2 Stochastic self-consistent harmonic approximation (SSCHA)
-In general, the free-energy of a ionic Hamiltonian $H=T+V$ is given by
-$F_{H}=\mathrm{tr}(\rho_{H}H)+\frac{1}{\beta}\mathrm{tr}(\rho_{H}\mathrm{ln}\rho_{H})$ (A2)
-where $\rho_{H}$ is the density matrix. In SSCHA, the density matrix is restricted to a trial harmonic Hamiltonian $\mathcal{H}=T+\mathcal{V}$, and the free energy is given by
-$\mathcal{F}_{H}(\mathcal{H})=\mathrm{tr}(\rho_{\mathcal{H}}H)+\frac{1}{\beta}\mathrm{tr}(\rho_{\mathcal{H}}\mathrm{ln}\rho_{\mathcal{H}})$ (A3)
-Due to a variational principle, $F_{H}\leq\mathcal{F}_{H}(\mathcal{H})$.
-The parameters of the harmonic Hamiltonian $\mathcal{H}$ include the centroids of the atoms, which correspond to the lattice parameters $a_{i}$ of the QHA, but it also includes the interatomic force constant matrix, which contains the information of all the phonon frequencies and accounts for the true anharmonicities.
-Thus, SSCHA has additional parameters, and is expected in general to give better results than QHA.
-Also, QHA does not satisfy a variational principle.
-It can be shown that
-[…]
-where $\mathbf{R}$ indicates a general ionic configuration and $\rho_{\mathcal{H}}(\mathbf{R})=\langle\mathbf{R}|e^{-\beta\mathcal{H}}|\mathbf{R}\rangle/Z_{\mathcal{H}}$ is the probability density of finding the crystal in a generic configuration $\mathbf{R}$ ($Z_{\mathcal{H}}=\mathrm{tr}[e^{-\beta\mathcal{H}}]$ the partition function).
-By stochastically sampling configurations $\mathbf{R}$ for a given density matrix $\rho_{\mathcal{H}}$, and determining the gradient of the free energy with respect the harmonic parameters (centroids and force constant matrix elements), the free energy can be minimized self-consistently.
-We used 24000 configurations in all calculations.
-A detailed description of SSCHA can be found for example in Refs.
-Errea et al., 2014 or Monacelli et al., 2021.
-The logarithm term in Eq.
-A1 can be hard to converge for low frequencies.
-So rather than calculating $F_{\mathcal{H}}$ directly with the grid corresponding to the SC used in the minimization procedure, the force constant matrix can be Fourier interpolated in the usual way, and $F_{\mathcal{H}}$ can be calculated using a dense grid.
-This is the novel SSCHA interpolation method we mentioned in the main section, which was used with a $40\times 40$ in-plane grid to determine the barrier and also the out-of-plane TE.
-To properly determine the barrier, the interpolation scheme is used both at equilibrium and at SP, and then the interpolated free energies are subtracted.
-The SSCHA frequencies are in principle auxiliary frequencies that minimize the free energy.
-In a purely harmonic system, such frequencies correspond to the physical frequencies.
-But in general, they have to be corrected with the so-called “bubble correction” (and also an additional higher order term in very anharmonic cases)Monacelli et al.
-(2021).
-Later in Fig.
-A5, we show the temperature dependendence of both the auxiliary and physical frequencies.
-A2 Barrier
+A1 Barrier
 To determine the barrier, the free energy has to be determined at the equilibrium position and SP.
@@
 3(b).
-As mentioned in the main text, the mode perpendicular to the bond direction shown in Fig.
+The mode perpendicular to the bond direction shown in Fig.
 1 is stable in SP, while the shear mode along the reaction coordinate (using the language of chemical reactions) is unstable.
@@
 A1), which makes the convergence of the free energy hard, mostly due to the logarithm term in $F_{\mathcal{H}}$ in Eq.
-(A4).
+(4).
 Just determining the barrier with this equation and a $N\times N\times 1$ supercell, results in Fig.
 A2, which is not well converged.
-In order to get better convergence, we use the interpolation method we mentioned in Sec.
-A1.2, and obtain a much more stable barrier as a function of $1/N$ (see Fig.
+In order to get better convergence, we use the interpolation method mentioned in Sec.
+II.2, and obtain a much more stable barrier as a function of $1/N$ (see Fig.
 A3).
@@
 Figure A2: Barrier as a function of temperature, for different supercells, just using the $N\times N\times 1$ grid to determine the harmonic part of the free energy Eq.
-(A4).
+(4).
 The barrier changes significantly with $N$, which means the shear mode is overrepresented.
@@
 Figure A3: Barrier as a function of $1/N$, for different temperatures, interpolating the first term of Eq.
-A4.
+4.
 The barrier seems well converged at low temperatures and harder to converge at higher temperatures.
@@
 Figure A4: SSCHA frequencies around the shear modes, at the equilibrium position (blue) and SP (red).
-A3 Shear and breathing frequency
+A2 Shear and breathing frequency
 The SSCHA auxiliary and physical frequencies as a function of temperature can be observed in Fig.
@@
 Figure A7: Convergence of the auxiliary (dashed) and physical (full) breathing frequency. The physical frequency is well-converged at $N=10$.
-A4 Thermal expansion (TE)
+A3 Thermal expansion (TE)
+For the SSCHA out-of-plane TE calculations, we also used the interpolation scheme with a $40\times 40$ grid.
 The interlayer distance and in-plane lattice parameter obtained in bulk with SSCHA can be seen in Fig.
 A8.
-The fit corresponds to Eq.
-(6) of Ref.
-Zorzi and Perottoni, 2021, both for in and out-of-plane TE ($b$ can be $a$ or $c_{d}$),
-$b(T)=b_{0}\mathrm{exp}\left(\frac{X\theta}{e^{X/T}-1}\right),$ (A5)
-which depends on parameters $b_{0}$, $X$ and $\theta$.
+The derivative of these fits results in the SSCHA curves of Fig.
+5.
 In Fig.
@@
 In what follows, we describe in detail convergence studies we did for SSCHA, and also for QHA calculations.
-Figure A8: Out-of plane (a) and in-plane (b) SSCHA thermal expansion (blue dots). The fit corresponds to Eq. (A5).
+Figure A8: Out-of plane (a) and in-plane (b) SSCHA thermal expansion (blue dots). The fit corresponds to Eq. (5).
 Figure A9: SSCHA and QHA out-of-plane coefficient of TE for graphite and bilayer graphene.
 Figure A10: In-plane coefficient of TE using SSCHA and QHA, for monolayer, bilayer and bulk.
-A4.1 Bilayer
+A3.1 Bilayer
 QHA.
@@
 A14).
-In-plane, a larger $N$ of about 50 is needed to fully achieve convergence (A15).
+In-plane, a larger $N$ of about 50 is needed to fully achieve convergence (Fig. A15).
 […]
-On the other hand, $a(T)$ is already converged at about $N=8$ (Fig. A17).
-A4.2 Bulk
+On the other hand, $a(T)$ has less variation with temperature relative to QHA and is already converged at about $N=8$ (Fig. A17).
+A3.2 Bulk
 QHA. In bulk, there is virtually no difference for the out-of-plane TE when fixing $a$, or considering also the minimization of $a$ (Fig. A19).
 In-plane (Fig. A20), there is a small difference between both fixing or varying $a$, similar to the bilayer case.
-SSCHA. Here we also used the interpolation method ($40\times 40$ grid), to go beyond the $N=8$ supercell that was possible in the direct SSCHA calculation, resulting in the blue dots of Fig. A8 a.
+SSCHA. Here we also used the interpolation method ($40\times 40$ grid), to go beyond the $N=8$ supercell that was possible in the direct SSCHA calculation, resulting in the blue dots of Fig. A8 (a).
 In-plane, based on $N=8$ already giving converged results in bilayer (Fig.
@@
 A15), $N=8$ should give converged SSCHA results in bulk, which corresponds to the blue dots in Fig.
-A8 b.
+A8 (b).
 Figure A11: Interlayer distance $c_{d}$ as a functi on of temperature in bilayer graphene using QHA, fixing the value of $a$ (dashed-blue) and doing the full minimization (full-black).

### v3

@@
-Quantum effects and temperature dependence of the free energy barrier and shear frequency in bilayer graphene
+Free energy barrier and thermal-quantum behavior of sliding bilayer graphene
 Abstract
-Little is known about the quantum effects of friction, and first-principles studies are virtually non-existent.
-In multilayer graphene, the order of the stacking of the layers plays a crucial role in the electronic properties and the manifestation of superconductivity, and friction affects how layers slide to change the stacking.
-Here, we study the effects of friction in bilayer graphene by focusing on quantum fluctuations in the free energy barrier that layers have to overcome to slide with respect to each other.
-We evaluate the free energy at the barrier in bilayer graphene through the self-consistent harmonic approximation, accounting for quantum and thermal ionic fluctuations, and obtain a very large reduction of the barrier of more than 30% at low temperatures.
-There is little temperature dependence up to about 500 K, but it increases at higher temperatures.
-Our approach is general and paves the way for systematically accounting for nuclear quantum and thermal effects in free energy barriers of other macroscopic systems.
+In multilayer graphene, the stacking order of the layers plays a crucial role in the electronic properties and the manifestation of superconductivity.
+By applying shear stress, it is possible to induce sliding between different layers, altering the stacking order.
+Here, focusing on bilayer graphene, we analyze how ionic fluctuations alter the free energy barrier between different stacking equilibria.
+We calculate the free energy barrier through the state-of-the-art self-consistent harmonic approximation, which can be evaluated at unstable configurations.
+We find that above 100 K there is a large reduction of the barrier of more than 30% due to thermal vibrations, which significantly improves the agreement between previous first-principles theoretical work and experiments in a single graphite crystal.
+As the temperature increases, the barrier remains nearly constant up to around 500 K, with a more pronounced decrease only at higher temperatures.
+Our approach is general and paves the way for systematically accounting for thermal effects in free energy barriers of other macroscopic systems.
 I Introduction
 There has been intense research on graphene since its discovery in 2004, and also on multilayer graphene systems.
-Among many other interesting properties, flat bands are present in multilayer rhombohedral (ABC-stacked) graphene (RG)Koshino (2010), and in fact, superconductivity has been observed in RG in 2021Zhou et al.
+Among many other interesting properties, flat bands are present in multilayer rhombohedral (ABC-stacked) graphene (RG)Koshino (2010), and superconductivity has been observed in RG in 2021Zhou et al.
 (2021).
@@
 (2020); Bouhafs et al.
-(2020), and thus of friction.
+(2021).
 To move graphene layers with respect to each other, they have to overcome a free energy barrier that separates different minima.
@@
 The maximum shear-stress to “unlock” layers is correlated with the height of the barrier.
-Such barrier corresponds to the activation energy of a metastable state, where a graphene flake (from nanometers up to possibly micrometer scales) sitting on top of a large graphene layer is displaced relative to the equilibrium position.
-In clean samples with micrometer sized flakes however, different regions can have different stackings (AB and AC)Henck et al.
-(2018); Jiang et al.
-(2018), and a domain wall in the metastable state separates both regions.
-The displacement of layers in this case is more complex and is not considered in our work.
-Superlubricity is another phenomenon present in systems involving graphene layers or graphite that has attracted a lot of attention(Liu et al., 2014; Sinclair et al., 2018), in which the energy landscape corresponding to the relative displacement of graphene layers is also central.
+Such barrier corresponds to the activation energy of a metastable state, where a graphene flake sitting on top of a large graphene layer is displaced relative to the equilibrium position.
+Superlubricity is a another phenomenon where the energy landscape corresponding to the relative displacement of layers is also central.
+In this mechanism, which has attracted a lot attention(Liu et al., 2014; Sinclair et al., 2018), particularly in carbon-based systems like multilayer graphene and graphite, incommensurate surfaces (i.e.
+atomic lattices without matching periodicity) lead to extremely low friction.
 Yet another phenomena involving sliding surfaces is thermolubricity, which refers to the assistance from thermal excitations to overcome barriers.
@@
 (2018).
-A few works have studied how quantum effects, including zero-point renormalization (ZPR) and tunneling, affect the free-energy barrier of dissociative adsorptionMills and Jónsson (1994) and diffusionPoma et al.
+Despite the interest in sliding surfaces, there are no studies that account for both temperature and quantum effects of free-energy barriers in macroscopic systems.
+A few works have studied quantum effects in free-energy barriers, including zero-point renormalization and tunneling, but of systems involving hydrogen molecules, including dissociative adsorptionMills and Jónsson (1994) and diffusionPoma et al.
 (2012); McIntosh et al.
 (2013); Cendagorta et al.
-(2016) of hydrogen.
-However, there are no studies of the quantum effects and temperature dependence of free-energy barriers in macroscopic systems.
+(2016).
+The variation of the potential along a periodic crystal surface, combined with the attraction between layers, prevents free sliding between them.
+This phenomenon has been extensively studied in various systems, like graphiteDienwiebel et al.
+(2004), MoS2Li et al.
+(2016), silicon on goldLiu et al.
+(2015), among others.
+When using a tip with the goal of sliding layers, normal forces need not be large.
+For pressures much lower than 1GPa, the alteration of the free-energy landscape is minimal (see SI of Ref.
+Nery et al., 2020), so the relaxed barrier remains the dominant contribution that impedes sliding.
+In addition, some works on superlubricity have observed that friction has an extremely weak dependence with the load, attributed for exampleDienwiebel et al.
+(2004) to a contact area that remains nearly constant in the force range applied.
+Here, we focus on calculating the height of the barrier, which should be approximately proportional to the force –given by the maximum slope of the free energy profile– that is needed for a layer to move from one minima to another in commensurate systems.
 A quantity closely related to the free energy barrier is the shear frequency, corresponding to the mode in which layers move rigidly in-plane and out of phase around the equilibrium position.
@@
 Ref.
-Mittal et al., 2021 uses the quasi-harmonic approxmimation (QHA), which underestimates the out-of-plane TE at low temperaturesBailey and Yates (1970) and has a very rapid increase at high temperatures; in-plane, it also underestimates the TE at low temperaturesBailey and Yates (1970) and overestimates it at high temperaturesMorgan (1972).
+Mittal et al., 2021 uses the quasi-harmonic approximation (QHA), which underestimates the out-of-plane TE at low temperaturesBailey and Yates (1970) and has a very rapid increase at high temperatures; in-plane, it also underestimates the TE at low temperaturesBailey and Yates (1970) and overestimates it at high temperaturesMorgan (1972).
 Much better out-of-plane results were obtained by using path integral molecular dynamics (PIMD)Herrero and Ramírez (2020), but while the TE flattens at 500 K in graphite, experimental values continue increasingMorgan (1972); Marsden et al.
@@
 Herrero and Ramírez, 2020 vastly overestimate the TE.
-Motivated by the aforementioned phenomena, especially by the lack of understanding of quantum effects in friction, we focus on studying the ZPR and temperature effects of the free-energy barrier.
+Motivated by the aforementioned phenomena, especially by the possibility of facilitating sliding through temperature, we focus on thermal effects on the free-energy barrier.
 This is a very challenging problem due to the difficulty of determining the free energy in a non-equilibrium position.
-Here we tackle it using the state-of-the-art stochastic self-consistent harmonic approximation (SSCHA)Monacelli et al.
+Here we tackle it using the state-of-the-art stochastic self-consistent harmonic approximation (SSCHA)Errea et al.
+(2014); Monacelli et al.
 (2021).
-First, in Sec.II, we briefly introduce SSCHA and the QHA, and explain how is the barrier determined.
+First, in Sec.
+II, we briefly introduce SSCHA and the QHA, and explain how the barrier is determined.
 Then, in Sec.
@@
 Here, we briefly describe the QHA, SSCHA, and a new SSCHA interpolation method used to determine the barrier.
-We also introduce the equation that is used to fit SSCHA thermal expansion results.
 A more detailed description of SSCHA can be found in Refs.
@@
 II.1 Quasi-harmonic approximation (QHA)
-In the QHA, at each value of the lattice parameters $\{a_{i}\}$, the free energy $F(T,\{a_{i}\})=U(\{a_{i}\})-TS(T,\{a_{i}\})$ is given by the standard harmonic expression. That is,
-$\begin{split}F(T,\{a_{i}\})=&U_{\mathrm{lat}}(\{a_{i}\})+\frac{1}{N}\sum_{\mathbf{q}\nu}\frac{1}{2}\hbar\omega_{\mathbf{q}\nu}(\{a_{i}\})+\\ &+\frac{1}{N}\sum_{\mathbf{q}\nu}kT\mathrm{ln}\left(1-e^{-\hbar\omega_{\mathbf{q}\nu}(\{a_{i}\})/kT}\right).\\ \end{split}$ (1)
+In the QHA, at each value of the lattice parameters $\{a_{i}\}$, the free energy $F(T,\{a_{i}\})=U(\{a_{i}\})-TS(T,\{a_{i}\})$, where $U$ is the internal energy, $S$ the entropy and $T$ the temperature, is given by the standard harmonic expression.
+That is,
+$\begin{split}F(T,\{a_{i}\})=&U_{\mathrm{lat}}(\{a_{i}\})+\frac{1}{N_{q}}\sum_{\mathbf{q}\nu}\frac{1}{2}\hbar\omega_{\mathbf{q}\nu}(\{a_{i}\})+\\ &+\frac{1}{N_{q}}\sum_{\mathbf{q}\nu}kT\mathrm{ln}\left(1-e^{-\hbar\omega_{\mathbf{q}\nu}(\{a_{i}\})/kT}\right).\\ \end{split}$ (1)
+where $U_{\mathrm{lat}}\{a_{i}\}$ is the internal energy of the lattice for a given value of the lattice parameters, $N_{q}$ the number of phonon wavevectors $\mathbf{q}$, $\nu$ the mode, and $\omega_{\mathbf{q}\nu}$ the corresponding phonon frequency.
 The value of the lattice parameters at a given temperature, $a_{i}(T)$, is determined by minimizing the free energy.
 Using an interatomic potential, the phonon frequencies at each value of the lattice parameters can be determined with little computational effort, and thus the QHA offers a quick method to determine the lattice expansion.
-Using first-principles calculations, phonons can be determined using density functional perturbation theory (DFPT), which uses a primitive cell and is much faster than frozen phonon (finite differences) calculations in supercells.
+Using first-principles calculations, phonons can be determined using density functional perturbation theory, which uses a primitive cell and is much faster than frozen phonon (finite differences) calculations in supercells.
 However, it does not account for the “true anharmonicity”Allen (2019), present at fixed values of the parameters when varying temperature.
 For example, it gives an increasing as opposed to decreasing temperature dependence of the G mode in graphene, due to the negative TE of the lattice parameter.
-Also, if the TE is negative and the lattice parameter reduces below the classical lattice value, as in the case of monolayer graphene at about 400 K (see later Fig.
-A10), then an acoustic mode becomes unstable and the QHA is ill-defined.
+Also, if the TE is negative and the lattice parameter reduces below the classical lattice value, as in the case of monolayer graphene at about 400 K (see the Supplementary Information (SI), Fig.
+S11), then an acoustic mode becomes unstable and the QHA is ill-defined.
 These issues are not present in SSCHA.
@@
 By stochastically sampling configurations $\mathbf{R}$ for a given density matrix $\rho_{\mathcal{H}}$, and determining the gradient of the free energy with respect the harmonic parameters (centroids and force constant matrix elements), the free energy can be minimized self-consistently.
+SSCHA is thus a nonperturbative method, as it does not rely on a perturbative expansion: it explores configurations with amplitudes that depend on temperature, without being limited to small displacements.
 Thanks to the variational ansatz, the SSCHA quantum free energy is rigorously defined, even in out-of-equilibrium and unstable configurations like SP.
-Notably, on the other hand, the QHA cannot determine the barrier, since the harmonic frequencies are not well defined at SP.
+In contrast, the QHA cannot determine the barrier, since the harmonic frequencies are not well defined at SP.
 The logarithm term in the harmonic term Eq.
@@
 (4) allows to interpolate the auxiliary harmonic term up to a very fine $q$-mesh, which is fundamental to avoid oversampling the energy of the long-ranged shear modes, while the anharmonic contribution can be properly converged already with small simulation cells.
-So rather than calculating $F_{\mathcal{H}}$ directly with the grid corresponding to the SC used in the minimization procedure, the force constant matrix can be Fourier interpolated in the usual way, and $F_{\mathcal{H}}$ can be calculated using a dense mesh.
+So rather than calculating $F_{\mathcal{H}}$ directly with the grid corresponding to the SC used in the minimization procedure, the force constant matrix can be Fourier interpolated in the usual way (building the force constant in real space from the existing $q$-grid, and then Fourier transforming back to any desired value of $\mathbf{q}$), and $F_{\mathcal{H}}$ can be calculated using a dense mesh.
 This procedure is followed both at SP and the equilibrium position, and the barrier is obtained by taking the difference of the corresponding free energies.
-More technical details are provided in the Appendix, Sec.
-A1.
+More technical details are provided in the SI, Sec.
+S1.
 The SSCHA frequencies are in principle auxiliary frequencies that minimize the free energy.
@@
 III.2 to obtain the temperature dependence of the shear frequency.
-Convergence with the supercell size of both the physical and auxiliary frequencies can be seen in the Appendix, Fig.
-A5.
-II.3 Thermal expansion (TE)
-As opposed to the barrier, the thermal expansion (TE) can also be calculated using the QHA, which although is much faster, is also in principle less accurate.
-With both QHA and SSCHA we determine the in-plane and out-of-plane lattice parameters as a function of temperature, $a(T)$ and $c(T)$, to then determine the in-plane and out-of-plane coefficients of TE, $\alpha_{\mathrm{a}}(T)=\tfrac{1}{a}\tfrac{da}{dT}$ and $\alpha_{\mathrm{c}}(T)=\tfrac{1}{c}\tfrac{dc}{dT}$.
-In the QHA case, the TE can calculated quickly and accurately for many temperatures, so we just use finite differences.
-For SSCHA, to reduce numerical error, we use Eq.
-(6) of Ref.
-Zorzi and Perottoni, 2021,
-$b(T)=b_{0}\mathrm{exp}\left(\frac{X\theta}{e^{X/T}-1}\right),$ (5)
-which depends on parameters $b_{0}$, $X$ and $\theta$, to fit $a(T)$ and $c_{d}(T)$ ($b$ can be $a$ or $c_{d}$). Then we determine the coefficients of TE from the derivative of the fit.
+Convergence with the supercell size of both the physical and auxiliary frequencies can be seen in the SI, Fig.
+S6.
 III Results
@@
 The energy profile of a layer moving with respect to the other one along the bond direction is given by Fig.
-1, which gives good reults when comparing to theory and experiment (see Table 1).
+1, which gives good results when comparing to theory and experiment (see Table 1).
 Table 1: Comparison of the barrier $V_{\mathrm{SP}}$ and the shear frequency at fixed ionic positions (classical at $T=0$ K) between the interatomic potential we used in this workLibbi et al.
@@
 SP is a saddle point in the full energy landscape, so the shear mode in the perpendicular direction (along the direction of two nearby AA maxima) is stable.
-In order to determine the ZPR and temperature dependence of the barrier, we use SSCHA to determine the free energy at the equilibrium position and SP, using the interpolation method described in Sec.
+In order to determine temperature effects on the barrier, we use SSCHA to determine the free energy at the equilibrium position and SP, using the interpolation method described in Sec.
 II.2.
+The barrier as a function of temperature can be seen in blue in Fig.
+2.
 Figure 1: Energy profile of bilayer graphene when moving one layer relative to the other along the bond direction.
@@
 It is reduced from 1.24 meV/atom, the value at fixed nuclei (VSP in Fig.
-1), to about 0.8 meV/atom, corresponding to a large reduction of about 35% due to anharmonicities.
+1), to about 0.8 meV/atom, corresponding to a large reduction of about 35% due to phonons.
 The barrier does not change much up to about 500 K, but the value at 1000 K is about 20% smaller relative to the value at 100 K.
 The lines are a guide to the eye, as in the other figures of this work.
-[…]
-To confirm that quantum fluctuations are very minor down to 100 K, we performed SSCHA calculations in the classical limit $\hbar\omega/kT\rightarrow 0$ (which recovers the classical Boltzmann distribution for the distorted configurations) at 100 K and 300 K.
-We obtained comparable results, except for a small difference of less than 0.05 meV/atom at 100 K with respect to the quantum values.
-Much lower temperatures would have to be considered to recover the classical value at fixed ions.
-Thus, it is the anharmonic vibrations that play a decisive role in the ability of layers to slide with respect to each other.
+The barrier at 100 K is reduced by a substantial 35% with respect to the value at fixed nuclei.
+We expect intralayer modes to contribute little to the reduction of the barrier, since they should not change much when displacing the layers.
+On the other hand, modes involving the relative motion of the layers should be much more affected.
+Indeed, intralayer modes are virtually identical, so they do not contribute to a change in the barrier.
+Regarding the interlayer modes, the breathing mode changes little, while the shear modes do differ significantly (see Fig.
+S4).
+Thus, the shear modes are responsible for the large reduction of the barrier.
+This makes sense intuitively: modes that involve a relative displacement of the layers should facilitate sliding.
+In the fixed ion calculation, the system is exactly at SP, but in reality, the system is also exploring surrounding configurations, which have a lower energy.
+Remarkably, our results show improved agreement with experiments.
 In a previous workNery et al.
 (2020), using density functional theory with an LDA functional (see Table 1), some of us obtained that the maximum shear stress in simulated stick-slip motion is 0.2 GPa, and we mentioned a good agreement with experimental values 0.14 GPa (Ref.
-Liu et al., 2012) and 0.1 GPa (Ref.
+Liu et al., 2012) –a benchmark for defect free single crystal graphiteBlomquist et al.
+(2019)– and 0.1 GPa (Ref.
 Dienwiebel et al., 2004).
-Assuming the maximum shear stress decreases by 35% just like the barrier, we obtain 0.13 GPa, in excellent agreement with experiments.
-Other systems involving hydrogen diffusion across a free energy barrier do present significant differences between classical and quantum free energy barriers due to the large quantum fluctuations of hydrogen.
-Such differences have been observed at 100 K (Ref.
-Mills and Jónsson, 1994), 250 K (Ref.
-McIntosh et al., 2013), and even higher temperaturesPoma et al.
-(2012), although other worksCendagorta et al.
-(2016) obtain very similar barriers at already 200 K.
+Assuming that the maximum shear stress decreases by 35% in line with the barrier, we now obtain 0.13 GPa, in excellent agreement with experimentsDienwiebel et al.
+(2004); Liu et al.
+(2012).
+[…]
+The size of the graphene flakes considered in our work in principle correspond to nanometers up to possibly micrometer scales.
+In clean samples with micrometer sized flakes, different regions can have different stackings (AB and AC)Henck et al.
+(2018); Jiang et al.
+(2018), and a domain wall in the metastable state separates both regions.
+The displacement of layers in this case is more complex and is not considered in our work.
+We expect shear stress to be lower in this case, since the metastable state is already present.
+Defects and stacking faults also lead to lower friction due to incommensurabilities, so the shear stress of 0.14 GPa corresponds to the maximum expected valueBlomquist et al.
+(2019).
+In other materials, the anharmonicities associated with the unstable modes at a saddle point should also contribute negatively to the free energy.
+However, the overall change of the barrier will depend on differences in the entire SSCHA phonon spectra at the minima and saddle point.
 Regarding the evolution of the barrier as the temperature increases, there is a weak dependence up to about 500 K, since the barrier remains close to around 0.8 meV/atom.
 However, when temperature increases up to 1000 K, the barrier is further reduced by 20% relative to the value at 100 K.
-Thus, high temperatures should help to reduce (static) friction between layers.
+Thus, high temperatures should facilitate sliding between layers.
 In particular, it could aid the production of ABC-stacked multilayer graphene from its more common AB-stacked version through shear stress, as mentioned in the Introduction.
@@
 In the QHA, the TE is obtained by determining the lattice parameters that minimize the free energy at each temperature.
-The third out of phase mode of the layers is the breathing mode, in the out-of-plane direction. A discussion of its temperature dependence can be found in the Appendix.
+The third out of phase mode of the layers is the breathing mode, in the out-of-plane direction. A discussion of its temperature dependence can be found in the SI.
 The temperature dependence of the shear frequency can be seen in Fig.
@@
 The quantum effects evaluated within the SSCHA harden the shear frequency energy at low temperatures compared with the classical value at 0 K (dashed line).
-On the other hand, the ZPR in QHA softens the shear frequency.
+On the other hand, the zero-point renormalization in QHA softens the shear frequency.
 The temperature dependence of the shear mode is similar to that of the barrier, varying by about 20% from 0 to 1000 K, and having a larger dependence at temperatures above 500 K.
@@
 The breathing mode, on the other hand, is flat, so nearby sampling points are not necessary to converge the free energy.
+Figure 4: Shear frequency as a function of temperature using SSCHA (blue circles) and QHA (red crosses).
+There is no experimental data for the temperature dependence of the shear frequency as a function of temperature for bilayer graphene.
+Experimental values correspond to bulk (blue crosses) and folded 2+2 graphene (orange circles), which are corrected by the factor of a nearest neighbor modelHe et al.
+(2016) to compare to the bilayer case (and rigidly shifted by -5 cm-1 to facilitate the visual comparison).
+The agreement with the temperature dependence of the bulk values is good.
 Regarding the experimental data, in Ref.
@@
 Cuscó et al., 2019, which show a similar temperature dependence (adjusting the bulk value by a $\sqrt{2}$ factor) of roughly -0.5 cm-1/100 K.
-Figure 4: Shear frequency as a function of temperature using SSCHA (blue circles) and QHA (red crosses).
-There is no experimental data for the temperature dependence of the shear frequency as a function of temperature for bilayer graphene.
-Experimental values correspond to bulk (blue crosses) and folded 2+2 graphene (orange circles), which are corrected by the factor of a nearest neighbor modelHe et al.
-(2016) to compare to the bilayer case (and rigidly shifted by -5 cm-1 to facilitate the visual comparison).
-The agreement with the temperature dependence of the bulk values is good.
 III.3 Thermal expansion
+Figure 5: (a) Graphite out-of-plane coefficient of TE as a function of temperature using SSCHA and QHA.
+PIMD calculationsHerrero and Ramírez (2020) and experimental data Bailey and Yates (1970) are also included.
+PIMD overall agrees well, while QHA and SSCHA (with the potential of Ref.
+Wen et al., 2018) underestimates the TE.
+(b) Graphite in-plane coefficient of TE using SSCHA and QHA.
+QHA again underestimates the dependence at low temperatures, and also at high temperatures.
+On the other hand, PIMD gives a dependence that is surprisingly very far from experimental values.
+SSCHA gives excellent results both at low and high temperatures, considering also the disagreement within experimental data.
 To further check the adequacy of our method (the use of SSCHA with the in-planeLibbi et al.
@@
 (2018) interatomic potentials), we look at the TE in the previous calculations and also in bulk graphite, and compare to experiments.
+The TE coefficients are given by $\alpha_{\mathrm{a}}(T)=\tfrac{1}{a}\tfrac{da}{dT}$ and $\alpha_{\mathrm{c}}(T)=\tfrac{1}{c}\tfrac{dc}{dT}$, with $a$ and $c$ the in-plane and out-of-plane lattice parameters, respectively.
 In Fig.
@@
 (2018) and PIMD calculations Herrero and Ramírez (2020).
-SSCHA curves result from fitting Eq.
-(5) to $a(T)$ and $c(T)$ (see Fig.
-A8) and then taking the derivative.
-SSCHA results underestimate the thermal expansion.
+SSCHA results underestimate the TE.
 PIMD results are overall very good, although they also give values somewhat lower than experimentsBailey and Yates (1970); Marsden et al.
@@
 QHA works well at low temperatures but departs from experiments at higher temperatures.
+We conducted a careful convergence of these results, reported in the SI, Sec.
+S3.
 PIMD calculations of Ref.
@@
 It would be interesting in future work to compare SSCHA and PIMD using the same potentials.
-Figure 5: (a) Graphite out-of-plane coefficient of thermal expansion as a function of temperature using SSCHA and QHA.
-PIMD calculationsHerrero and Ramírez (2020) and experimental data Bailey and Yates (1970) are also included.
-PIMD overall agrees well, while QHA and SSCHA (with the potential of Ref.
-Wen et al., 2018) underestimates the TE.
-(b) Graphite in-plane coefficient of thermal expansion using SSCHA and QHA.
-QHA again underestimates the dependence at low temperatures, and also at high temperatures.
-On the other hand, PIMD gives a dependence that is surprisingly very far from experimental values.
-SSCHA gives excellent results both at low and high temperatures, considering also the disagreement within experimental data.
 IV Conclusions
-To conclude, we demonstrated that anharmonicities are fundamental to determining properties of layered materials that account for friction.
-Using the self-consistent harmonic approximation, which can calculate the quantum free energy at unstable configurations, and a novel interpolation method, we unveiled the suppression of the barrier separating stable configurations in bilayer graphene.
-Anharmonicities reduce the free energy barrier by 35%, and gets further reduced at high temperatures.
-On the other hand, quantum fluctuations are not important down to 100 K.
+To conclude, we demonstrated that ionic vibrations are fundamental to determining properties of layered materials that depend on free energy barriers.
+Using the self-consistent harmonic approximation, which can calculate the free energy at unstable configurations, and a novel interpolation method, we unveiled the suppression of the barrier separating stable configurations in bilayer graphene.
+Phonon vibrations reduce the free energy barrier by a significant 35%, and gets further suppressed at high temperatures, so barriers at fixed ions cannot be trusted.
+Now, correcting the shear stress of 0.2 GPa reported in our previous work, we arrive at a value of 0.13 GPa, in excellent agreement with experiments.
 The shear modes, which occupy a small region of phase space, have to be adequately sampled to correctly determine the free energy barrier as well as the thermal expansion.
@@
 101151380.
-Appendix
-A1 Barrier
+Supplementary Information
+S1 Barrier
 To determine the barrier, the free energy has to be determined at the equilibrium position and SP.
@@
 This is more problematic at SP, where the SSCHA frequency corresponding to she unstable mode takes very small values (see Fig.
-A1), which makes the convergence of the free energy hard, mostly due to the logarithm term in $F_{\mathcal{H}}$ in Eq.
+S1), which makes the convergence of the free energy hard, mostly due to the logarithm term in $F_{\mathcal{H}}$ in Eq.
 (4).
 Just determining the barrier with this equation and a $N\times N\times 1$ supercell, results in Fig.
-A2, which is not well converged.
-In order to get better convergence, we use the interpolation method mentioned in Sec.
+S2, which is not well converged.
+All SSCHA calculations used 24000 configurations, both for the barrier and the TE.
+In order to get better convergence with $N$, we use the interpolation method mentioned in Sec.
 II.2, and obtain a much more stable barrier as a function of $1/N$ (see Fig.
-A3).
+S3).
 The $40\times 40$ interpolation grid used for $F_{\mathcal{H}}$ corresponds to a supercell of 6400 atoms.
@@
 To better understand the reduction in the barrier, we show in Fig.
-A4 the SSCHA frequencies (around the shear modes) at equilibrium and SP at $T=100$ K.
+S4 the SSCHA frequencies (around the shear modes) at equilibrium and SP at $T=100$ K.
 The SSCHA frequencies at equilibrium are similar to the standard harmonic values, as shown in the main text.
 The SSCHA shear frequencies at SP instead change significantly.
-The shear mode that looks like an acoustic mode takes a value of about 2 cm-1 at $\Gamma$ (see also Fig.
-A1).
+The shear mode that looks like an acoustic mode takes a value close to 2 cm-1 at $\Gamma$ (see also Fig.
+S1).
 The rest of the SSCHA phonon dispersion presents small differences between equilibrium and SP frequencies.
-Figure A1: Unstable shear frequency at SP as a function of $1/N$, for different temperatures. The fact that it does not stabilize shows that the shear mode is present in a small region of the BZ.
-Figure A2: Barrier as a function of temperature, for different supercells, just using the $N\times N\times 1$ grid to determine the harmonic part of the free energy Eq.
+In addition, we performed calculations in the classical limit ($\hbar\omega/kT\rightarrow 0$) by increasing the carbon atomic masses by a factor of 600.
+This reduces the Debye temperature below 100 K, ensuring that all modes are thermally populated and recovering the classical Boltzmann distribution for the distorted configurations.
+As we can see in Fig.
+S5, the standard quantum SSCHA calculation and its classical limit coincide, showing quantum effects are neglegible above 100 K.
+Figure S1: Unstable shear frequency at SP as a function of $1/N$, for different temperatures. The fact that it does not stabilize shows that the shear mode is present in a small region of the BZ.
+Figure S2: Barrier as a function of temperature, for different supercells, just using the $N\times N\times 1$ grid to determine the harmonic part of the free energy Eq.
 (4).
@@
 See Fig.
-A3.
-Figure A3: Barrier as a function of $1/N$, for different temperatures, interpolating the first term of Eq.
+S3.
+Figure S3: Barrier as a function of $1/N$, for different temperatures, interpolating the first term of Eq.
 4.
@@
 Values are likely converged within less than 0.05 meV/atom.
-Figure A4: SSCHA frequencies around the shear modes, at the equilibrium position (blue) and SP (red).
-A2 Shear and breathing frequency
+Figure S4: SSCHA frequencies around the shear modes, at the equilibrium position (blue) and SP (red).
+Figure S5: Comparison between the standard quantum SSCHA results and its classical limit.
+S2 Shear and breathing frequency
 The SSCHA auxiliary and physical frequencies as a function of temperature can be observed in Fig.
-A5.
+S6.
 The frequencies are not fully converged, and the converged value might be similar to the QHA.
@@
 Its temperature dependence can be seen in Fig.
-A6.
+S7.
 It decreases more at low temperatures compared to the shear mode and barrier.
@@
 The auxiliary frequency for $N=8$ and 10 is almost the same (Fig.
-A7), so the physical values are also converged (within 0.5 cm-1 at 1000 K, and less for lower temperatures).
-Figure A5: Shear frequency as a function of temperature, both for the physical frequency (full lines) and the auxiliary one (dashed).
+S8), so the physical values are also converged (within 0.5 cm-1 at 1000 K, and less for lower temperatures).
+Figure S6: Shear frequency as a function of temperature, both for the physical frequency (full lines) and the auxiliary one (dashed).
 The experimental value is shifted and included twice, to facilitate the comparison of the temperature dependence with the QHA and SSCHA curves.
 The agreement of the experimental bulk value (adjusted with the nearest neighbor model to the bilayer case) with QHA is excellent, and with SSCHA is very good as well.
-Figure A6: Breathing frequency as a function of temperature using SSCHA (black circles) and QHA (red crosses).
+Figure S7: Breathing frequency as a function of temperature using SSCHA (black circles) and QHA (red crosses).
 The SSCHA calculations are well converged and serve as a prediction, since there is no experimental data in the literature.
-Figure A7: Convergence of the auxiliary (dashed) and physical (full) breathing frequency. The physical frequency is well-converged at $N=10$.
-A3 Thermal expansion (TE)
+Figure S8: Convergence of the auxiliary (dashed) and physical (full) breathing frequency. The physical frequency is well-converged at $N=10$.
+S3 Thermal expansion (TE)
 For the SSCHA out-of-plane TE calculations, we also used the interpolation scheme with a $40\times 40$ grid.
 The interlayer distance and in-plane lattice parameter obtained in bulk with SSCHA can be seen in Fig.
-A8.
-The derivative of these fits results in the SSCHA curves of Fig.
-5.
+S9.
+The fits are used to reduce numerical error, and correspond Eq.
+(6) of Ref.
+Zorzi and Perottoni, 2021,
+$b(T)=b_{0}\mathrm{exp}\left(\frac{X\theta}{e^{X/T}-1}\right),$ (S1)
+which depends on parameters $b_{0}$, $X$ and $\theta$ ($b$ can be $a$ or $c_{d}$).
+The derivative of the fits give the coefficients of TE.
+In the QHA case, the TE can be calculated quickly and accurately for many temperatures, so we just use finite differences.
 In Fig.
-A9 we compare SSCHA and QHA for bilayer graphene and bulk.
+S10 we compare SSCHA and QHA for bilayer graphene and bulk.
 The TE for both SSCHA and QHA bulk is lower relative to bilayer.
 In-plane, Fig.
-A10, we can also see a clear reduction of the TE coefficient from monolayer to bulk.
+S11, we can also see a clear reduction of the TE coefficient from monolayer to bulk.
 In what follows, we describe in detail convergence studies we did for SSCHA, and also for QHA calculations.
-Figure A8: Out-of plane (a) and in-plane (b) SSCHA thermal expansion (blue dots). The fit corresponds to Eq. (5).
-Figure A9: SSCHA and QHA out-of-plane coefficient of TE for graphite and bilayer graphene.
-Figure A10: In-plane coefficient of TE using SSCHA and QHA, for monolayer, bilayer and bulk.
-A3.1 Bilayer
+Figure S9: Out-of plane (a) and in-plane (b) SSCHA thermal expansion (blue dots). The fit corresponds to Eq. (S1).
+Figure S10: SSCHA and QHA out-of-plane coefficient of TE for graphite and bilayer graphene.
+Figure S11: In-plane coefficient of TE using SSCHA and QHA, for monolayer, bilayer and bulk.
+S3.1 Bilayer
 QHA.
@@
 In Fig.
-A11 we can see $c_{d}(T)$ both for fixed $a$, and for the full minimization.
+S12 we can see $c_{d}(T)$ both for fixed $a$, and for the full minimization.
 The full minimization has a slightly larger dependence with temperature, but fixing $a$ is a good approximation.
 Fig.
-A12 is analogous for $a(T)$, with a fixed value $c_{d}=3.5$.
+S13 is analogous for $a(T)$, with a fixed value $c_{d}=3.5$.
 The curves are very similar, except for a small difference at high temperatures (of about 0.0005 Å at 1000 K).
 We also studied convergence with the size of the grid (Fig.
-A13).
+S14).
 At $N=12$, $c_{d}(T)$ is well converged (within 0.005 Å at 1000 K), although $N\sim 20$ to achieve full convergence.
-The shear mode increases thermal expansion because when layers are displaced relative to each other, the interlayer distance is larger (see Fig.
-A14).
-In-plane, a larger $N$ of about 50 is needed to fully achieve convergence (Fig. A15).
+The shear mode increases TE because when layers are displaced relative to each other, the interlayer distance is larger (see Fig.
+S15).
+The monotonous behavior indicates that the shear modes are overrepresented for coarser grids, and become better sampled with denser grids (having wavevectors $\mathbf{q}$ closer to $\Gamma$).
+In-plane, a larger $N$ of about 50 is needed to fully achieve convergence (Fig.
+S16).
+Again, monotonous convergence suggests inadequate sampling close to $\Gamma$.
+Since the convergence in bulk is much faster (see below Fig.
+S24), we attribute the slow convergence to the flexural mode, which becomes linear (an acoustic mode) and easier to sample in bulk.
 […]
-On the other hand, $a(T)$ has less variation with temperature relative to QHA and is already converged at about $N=8$ (Fig. A17).
-A3.2 Bulk
-QHA. In bulk, there is virtually no difference for the out-of-plane TE when fixing $a$, or considering also the minimization of $a$ (Fig. A19).
-In-plane (Fig. A20), there is a small difference between both fixing or varying $a$, similar to the bilayer case.
-SSCHA. Here we also used the interpolation method ($40\times 40$ grid), to go beyond the $N=8$ supercell that was possible in the direct SSCHA calculation, resulting in the blue dots of Fig. A8 (a).
+On the other hand, $a(T)$ has less variation with temperature relative to QHA and is already converged at about $N=8$ (Fig. S18).
+S3.2 Bulk
+QHA.
+In bulk, there is virtually no difference for the out-of-plane TE when fixing $a$, or considering also the minimization of $a$ (Fig.
+S20).
+In-plane (Fig.
+S21), there is a small difference between both fixing or varying $a$, similar to the bilayer case.
+Just like in the bilayer case, relative dense grids are needed to achieve out-of-plane convergence (Fig.
+S22), which is again likely due to the shear mode.
+Regarding the extension of the supercell in the out-of-plane direction, 2 unit cells (4 layers) give good results (Fig.
+S23).
+In-plane (Fig.
+S24), convergence is much faster than out-of-plane.
+SSCHA. Here we also used the interpolation method ($40\times 40$ grid), to go beyond the $N=8$ supercell that was possible in the direct SSCHA calculation, resulting in the blue dots of Fig. S9 (a).
 In-plane, based on $N=8$ already giving converged results in bilayer (Fig.
-A17), and also considering that the QHA convergence in bulk (Fig.
-A23) is much faster than in bilayer (Fig.
-A15), $N=8$ should give converged SSCHA results in bulk, which corresponds to the blue dots in Fig.
-A8 (b).
-Figure A11: Interlayer distance $c_{d}$ as a functi on of temperature in bilayer graphene using QHA, fixing the value of $a$ (dashed-blue) and doing the full minimization (full-black).
+S18), and also considering that the QHA convergence in bulk (Fig.
+S24) is much faster than in bilayer (Fig.
+S16), $N=8$ should give converged SSCHA results in bulk, which corresponds to the blue dots in Fig.
+S9 (b).
+Figure S12: Interlayer distance $c_{d}$ as a functi on of temperature in bilayer graphene using QHA, fixing the value of $a$ (dashed-blue) and doing the full minimization (full-black).
 In-plane TE has little effect on the out-of-plane TE.
-Figure A12: Lattice parameter $a$ as a function of temperature in bilayer graphene using QHA, fixing the value of $c_{d}$ (dashed-blue) and doing the full minimization (full-black).
+Figure S13: Lattice parameter $a$ as a function of temperature in bilayer graphene using QHA, fixing the value of $c_{d}$ (dashed-blue) and doing the full minimization (full-black).
 In-plane TE has little effect on the out-of-plane TE.
-Figure A13: Interlayer distance as a function of temperture in bilayer graphene, using QHA. At $N\sim 20$ $c_{d}$ is converged, although $N=12$ also gives a good result.
-Figure A14: Interlayer distance profile along the bond direction, when moving a graphene layer relative to another one (analogous to Fig.
+Figure S14: Interlayer distance as a function of temperture in bilayer graphene, using QHA. At $N\sim 20$ $c_{d}$ is converged, although $N=12$ also gives a good result.
+Figure S15: Interlayer distance profile along the bond direction, when moving a graphene layer relative to another one (analogous to Fig.
 1, but for distance instead of energy).
 The minimum distance is at the equilibrium positions.
-Figure A15: In-plane lattice parameter as a function of temperature in bilayer graphene for different $N$ using QHA. A large $N$ is needed to achieve convergence.
-Figure A16: Interlayer distance as a function of temperature in bilayer graphene for different $N$ using SSCHA.
+Figure S16: In-plane lattice parameter as a function of temperature in bilayer graphene for different $N$ using QHA. A large $N$ is needed to achieve convergence.
+Figure S17: Interlayer distance as a function of temperature in bilayer graphene for different $N$ using SSCHA.
 Convergence is similar to the QHA case, Fig.
-A13, where $N=12$ gives a good value of $c_{d}(T)$.
-Figure A17: In-plane lattice parameter as a function of temperature, for different $N$. At $N=8$ the curve is already well converged.
-Figure A18: Interlayer distance as a function of temperature in bilayer graphene, using SSCHA, at two different values of the in-plane lattice parameter.
+S14, where $N=12$ gives a good value of $c_{d}(T)$.
+Figure S18: In-plane lattice parameter as a function of temperature, for different $N$. At $N=8$ the curve is already well converged.
+Figure S19: Interlayer distance as a function of temperature in bilayer graphene, using SSCHA, at two different values of the in-plane lattice parameter.
 The curves are similar to the ones of Fig.
-A11.
+S12.
 In-plane TE has little effect on the out-of-plane TE.
-Figure A19: Interlayer distance as a function of temperature in bulk using QHA, fixing the value of $c_{d}$ (dashed-blue) and doing the full minimization (full-black).
+Figure S20: Interlayer distance as a function of temperature in bulk using QHA, fixing the value of $c_{d}$ (dashed-blue) and doing the full minimization (full-black).
 In-plane TE has barely no effect on the out-of-plane TE.
-Figure A20: In-plane lattice parameter as a function of temperature in bulk using QHA, fixing the value of $c_{d}$ (dashed-blue) and doing the full minimization (full-black).
+Figure S21: In-plane lattice parameter as a function of temperature in bulk using QHA, fixing the value of $c_{d}$ (dashed-blue) and doing the full minimization (full-black).
 The out-of-plane TE has a slightly larger but similar effect relative to bilayer on the in-plane TE.
-Figure A21: Interlayer distance as a function of temperature in graphite for different $N$ using QHA. Convergence is similar to the bilayer case.
-Figure A22: Interlayer distance as a function of temperature in graphite for different $N_{z}$. $N_{z}=2$ is the value used in our calculations.
-Figure A23: In-plane lattice parameter as a function of temperature in graphite using QHA, for different $N$. At $N=6$ convergence is already very good, as opposed to the bilayer case Fig. A16.
+Figure S22: Interlayer distance as a function of temperature in graphite for different $N$ using QHA. Convergence is similar to the bilayer case.
+Figure S23: Interlayer distance as a function of temperature in graphite for different $N_{z}$. $N_{z}=2$ is the value used in our calculations.
+Figure S24: In-plane lattice parameter as a function of temperature in graphite using QHA, for different $N$. At $N=6$ convergence is already very good, as opposed to the bilayer case Fig. S17.

### v4

@@
 There has been intense research on graphene since its discovery in 2004, and also on multilayer graphene systems.
-Among many other interesting properties, flat bands are present in multilayer rhombohedral (ABC-stacked) graphene (RG)Koshino (2010), and superconductivity has been observed in RG in 2021Zhou et al.
-(2021).
-Several works recently studied, both experimentally and theoretically, under which conditions RG can form Yankowitz et al.
-(2014); Geisenhof et al.
-(2019); Bouhafs et al.
-(2021); Gao et al.
-(2020); Li et al.
-(2020); Kerelsky et al.
-(2021).
-In particular, some of us unveiled how the stacking of graphene layers changes under shear stress Nery et al.
-(2020), and obtained that it produces many consecutive layers of RG.
-In fact, all samples with long-range RG that have been observed so far have been subject to some form of shear stress Lin et al.
-(2012); Henni et al.
-(2016); Henck et al.
-(2018); Yang et al.
-(2019); Shi et al.
-(2020); Bouhafs et al.
-(2021).
+Among many other interesting properties, flat bands are present in multilayer rhombohedral (ABC-stacked) graphene (RG)1, and superconductivity has been observed in RG in 20212.
+Several works recently studied, both experimentally and theoretically, under which conditions RG can form 3, 4, 5, 6, 7, 8, given that AB-stacked graphene is the more common occurring phase.
+In particular, some of us unveiled how the stacking of graphene layers changes under shear stress 9, and obtained that it produces many consecutive layers of RG.
+In fact, all samples with long-range RG that have been observed so far have been subject to some form of shear stress 10, 11, 12, 13, 14, 5.
 To move graphene layers with respect to each other, they have to overcome a free energy barrier that separates different minima.
-When doing so, they experiment “stick-slip” motion: they move gradually with increasing shear stress, followed by sudden rearrangements and decrease of the stressLiu et al.
-(2012).
+When doing so, they experiment “stick-slip” motion: they move gradually with increasing shear stress, followed by sudden rearrangements and decrease of the stress15.
 The maximum shear-stress to “unlock” layers is correlated with the height of the barrier.
@@
 Superlubricity is a another phenomenon where the energy landscape corresponding to the relative displacement of layers is also central.
-In this mechanism, which has attracted a lot attention(Liu et al., 2014; Sinclair et al., 2018), particularly in carbon-based systems like multilayer graphene and graphite, incommensurate surfaces (i.e.
+In this mechanism, which has attracted a lot attention16, 17, particularly in carbon-based systems like multilayer graphene and graphite, incommensurate surfaces (i.e.
 atomic lattices without matching periodicity) lead to extremely low friction.
 Yet another phenomena involving sliding surfaces is thermolubricity, which refers to the assistance from thermal excitations to overcome barriers.
-This term has been coined by some works that have studied the transition from stick-slip motion into stochastic fluctuations in friction at sufficiently high temperaturesBaykara et al.
-(2018).
+This term has been coined by some works that have studied the transition from stick-slip motion into stochastic fluctuations in friction at sufficiently high temperatures18.
 Despite the interest in sliding surfaces, there are no studies that account for both temperature and quantum effects of free-energy barriers in macroscopic systems.
-A few works have studied quantum effects in free-energy barriers, including zero-point renormalization and tunneling, but of systems involving hydrogen molecules, including dissociative adsorptionMills and Jónsson (1994) and diffusionPoma et al.
-(2012); McIntosh et al.
-(2013); Cendagorta et al.
-(2016).
+A few works have studied quantum effects in free-energy barriers, including zero-point renormalization and tunneling, but of systems involving hydrogen molecules, including dissociative adsorption19 and diffusion20, 21, 22.
 The variation of the potential along a periodic crystal surface, combined with the attraction between layers, prevents free sliding between them.
-This phenomenon has been extensively studied in various systems, like graphiteDienwiebel et al.
-(2004), MoS2Li et al.
-(2016), silicon on goldLiu et al.
-(2015), among others.
-When using a tip with the goal of sliding layers, normal forces need not be large.
+This phenomenon has been extensively studied in various systems, like graphite23, MoS224, silicon on gold25, among others.
+When using a tip with the goal of making layers slide, normal forces need not be large.
 For pressures much lower than 1GPa, the alteration of the free-energy landscape is minimal (see SI of Ref.
-Nery et al., 2020), so the relaxed barrier remains the dominant contribution that impedes sliding.
-In addition, some works on superlubricity have observed that friction has an extremely weak dependence with the load, attributed for exampleDienwiebel et al.
-(2004) to a contact area that remains nearly constant in the force range applied.
+9), so the relaxed barrier remains the dominant contribution that impedes sliding.
+In addition, some works on superlubricity have observed that friction has an extremely weak dependence with the load, attributed for example23 to a contact area that remains nearly constant in the force range applied.
 Here, we focus on calculating the height of the barrier, which should be approximately proportional to the force –given by the maximum slope of the free energy profile– that is needed for a layer to move from one minima to another in commensurate systems.
@@
 The free energy barrier separates two equivalent minima, so the barrier and shear frequency are expected to have a similar temperature dependence (a lower curvature at the minima corresponds to a lower shear frequency and a lower free energy, at least close to the minima).
-The value of the shear frequency in graphite is a relatively well-established quantity, with experimental valuesLebedeva et al.
-(2017) oscillating between 42 - 45 cm-1, using for example inelastic X-ray scatteringMohr et al.
-(2007), neutron-coherent inelastic scatteringNicklow et al.
-(1972), Raman spectraTan et al.
-(2012); Hanfland et al.
-(1989), and coherent phonon spectroscopy Boschetto et al.
-(2013).
+The value of the shear frequency in graphite is a relatively well-established quantity, with experimental values26 oscillating between 42 - 45 cm-1, using for example inelastic X-ray scattering27, neutron-coherent inelastic scattering28, Raman spectra29, 30, and coherent phonon spectroscopy 31.
 In bilayer graphene, there are some experiments with values around 30 cm-1; for example, 28 $\pm$ 3 cm-1 in Refs.
-S Amelinckx, 1965; Boschetto et al., 2013, and 32 cm-1 in Ref.
-Tan et al., 2012.
+32, 31, and 32 cm-1 in Ref.
+29.
 The temperature dependence of the shear modes, however, has been barely studied.
-Few experiments measure the temperature dependence of the shear mode, using Raman scattering in folded 2+2 (4 layer) grapheneCong and Yu (2014), or femtosecond pump-probe spectroscopy in graphiteMishina et al.
-(2000).
+Few experiments measure the temperature dependence of the shear mode, using Raman scattering in folded 2+2 (4 layer) graphene33, or femtosecond pump-probe spectroscopy in graphite34.
 Another aspect of multilayer graphene and graphite that still needs further investigation and that serves as a consistency check on the adequacy of our approach is thermal expansion (TE).
 Ref.
-McQuade et al., 2021 measured it in bilayer and trilayer graphene for the first time.
+35 measured it in bilayer and trilayer graphene for the first time.
 Ref.
-Mittal et al., 2021 uses the quasi-harmonic approximation (QHA), which underestimates the out-of-plane TE at low temperaturesBailey and Yates (1970) and has a very rapid increase at high temperatures; in-plane, it also underestimates the TE at low temperaturesBailey and Yates (1970) and overestimates it at high temperaturesMorgan (1972).
-Much better out-of-plane results were obtained by using path integral molecular dynamics (PIMD)Herrero and Ramírez (2020), but while the TE flattens at 500 K in graphite, experimental values continue increasingMorgan (1972); Marsden et al.
-(2018).
+36 uses the quasi-harmonic approximation (QHA), which underestimates the out-of-plane TE at low temperatures37 and has a very rapid increase at high temperatures; in-plane, it also underestimates the TE at low temperatures37 and overestimates it at high temperatures38.
+Much better out-of-plane results were obtained by using path integral molecular dynamics (PIMD)39, but while the TE flattens at 500 K in graphite, experimental values continue increasing38, 40.
 However, in-plane PIMD results of Ref.
-Herrero and Ramírez, 2020 vastly overestimate the TE.
+39 vastly overestimate the TE.
 Motivated by the aforementioned phenomena, especially by the possibility of facilitating sliding through temperature, we focus on thermal effects on the free-energy barrier.
 This is a very challenging problem due to the difficulty of determining the free energy in a non-equilibrium position.
-Here we tackle it using the state-of-the-art stochastic self-consistent harmonic approximation (SSCHA)Errea et al.
-(2014); Monacelli et al.
-(2021).
+Here we tackle it using the state-of-the-art stochastic self-consistent harmonic approximation (SSCHA)41, 42.
 First, in Sec.
@@
 II Theoretical and Computational Framework
-Here, we briefly describe the QHA, SSCHA, and a new SSCHA interpolation method used to determine the barrier.
-A more detailed description of SSCHA can be found in Refs.
-Errea et al., 2014 or Monacelli et al., 2021.
+Here, we briefly describe the QHA, SSCHA, and a new SSCHA interpolation method used to determine the barrier. A more detailed description of SSCHA can be found in Refs. 41 or 42.
 II.1 Quasi-harmonic approximation (QHA)
@@
 That is,
-$\begin{split}F(T,\{a_{i}\})=&U_{\mathrm{lat}}(\{a_{i}\})+\frac{1}{N_{q}}\sum_{\mathbf{q}\nu}\frac{1}{2}\hbar\omega_{\mathbf{q}\nu}(\{a_{i}\})+\\ &+\frac{1}{N_{q}}\sum_{\mathbf{q}\nu}kT\mathrm{ln}\left(1-e^{-\hbar\omega_{\mathbf{q}\nu}(\{a_{i}\})/kT}\right).\\ \end{split}$ (1)
+[…]
 where $U_{\mathrm{lat}}\{a_{i}\}$ is the internal energy of the lattice for a given value of the lattice parameters, $N_{q}$ the number of phonon wavevectors $\mathbf{q}$, $\nu$ the mode, and $\omega_{\mathbf{q}\nu}$ the corresponding phonon frequency.
@@
 Using first-principles calculations, phonons can be determined using density functional perturbation theory, which uses a primitive cell and is much faster than frozen phonon (finite differences) calculations in supercells.
-However, it does not account for the “true anharmonicity”Allen (2019), present at fixed values of the parameters when varying temperature.
+However, it does not account for the “true anharmonicity”43, present at fixed values of the parameters when varying temperature.
 For example, it gives an increasing as opposed to decreasing temperature dependence of the G mode in graphene, due to the negative TE of the lattice parameter.
 Also, if the TE is negative and the lattice parameter reduces below the classical lattice value, as in the case of monolayer graphene at about 400 K (see the Supplementary Information (SI), Fig.
-S11), then an acoustic mode becomes unstable and the QHA is ill-defined.
+S12), then an acoustic mode becomes unstable and the QHA is ill-defined.
 These issues are not present in SSCHA.
@@
 Due to a variational principle, $F_{H}\leq\mathcal{F}_{H}(\mathcal{H})$.
-The parameters of the harmonic Hamiltonian $\mathcal{H}$ include the centroids of the atoms, which correspond to the lattice parameters $a_{i}$ of the QHA, but it also includes the interatomic force constant matrix, which contains the information of all the phonon frequencies and accounts for the true anharmonicities.
+The parameters of the harmonic Hamiltonian $\mathcal{H}$ include the centroids of the atoms (their average positions), which correspond to the lattice parameters $a_{i}$ of the QHA, but it also includes the interatomic force constant matrix, which contains the information of all the phonon frequencies and accounts for the true anharmonicities.
 Thus, SSCHA has additional parameters, and is expected in general to give better results than QHA.
@@
 It can be shown that
-[…]
+$\mathcal{F}_{H}(\mathcal{H})=F_{\mathcal{H}}+\int d\mathbf{R}[V(\mathbf{R})-\mathcal{V}(\mathbf{R})]\rho_{\mathcal{H}}(\mathbf{R}),$ (4)
 where $\mathbf{R}$ indicates a general ionic configuration and $\rho_{\mathcal{H}}(\mathbf{R})=\langle\mathbf{R}|e^{-\beta\mathcal{H}}|\mathbf{R}\rangle/Z_{\mathcal{H}}$ is the probability density of finding the crystal in a generic configuration $\mathbf{R}$ ($Z_{\mathcal{H}}=\mathrm{tr}[e^{-\beta\mathcal{H}}]$ is the partition function).
@@
 In a purely harmonic system, such frequencies correspond to the physical frequencies.
-But in general, they have to be corrected with the so-called “bubble correction” (and also an additional higher order term in very anharmonic cases)Monacelli et al.
-(2021), which we do in Sec.
+But in general, they have to be corrected with the so-called “bubble correction” (and also an additional higher order term in very anharmonic cases)42, which we do in Sec.
 III.2 to obtain the temperature dependence of the shear frequency.
 Convergence with the supercell size of both the physical and auxiliary frequencies can be seen in the SI, Fig.
-S6.
+S7.
+Since the centroids of the atoms are also parameters in SSCHA, from the minimization of the free energy we additionally obtain the lattice parameters as a function of temperature, allowing us to study their thermal expansion.
+This is discussed in Sec.
+III C.
 III Results
@@
 We first focus on studying bilayer graphene, where the in-plane and (most importantly) out-of-plane interactions are given by the interatomic potential of Ref.
-Libbi et al., 2020 and Ref.
-Wen et al., 2018, respectively.
+44 and Ref.
+45, respectively.
 The out-of-plane interactions determine the barrier.
-The energy profile of a layer moving with respect to the other one along the bond direction is given by Fig.
-1, which gives good results when comparing to theory and experiment (see Table 1).
-Table 1: Comparison of the barrier $V_{\mathrm{SP}}$ and the shear frequency at fixed ionic positions (classical at $T=0$ K) between the interatomic potential we used in this workLibbi et al.
-(2020); Wen et al.
-(2018), LDA, and theoretical calculations and experiments from previous works.
-The values are a little bit lower but close to that of adiabatic-connection fluctuation-dissipation theorem within the random phase approximation (ACFDT-RPA)Zhou et al.
-(2015) and experimental dataLebedeva et al.
-(2017).
+The energy profile of a layer moving with respect to the other along the bond direction, shown in Fig.
+1, is obtained by fixing the in-plane positions of the layers and relaxing the interlayer distance until the out-of-plane components of the atomic forces vanish, and gives good agreement with theory and experiment (see Table 1).
+Table 1: Comparison of the barrier $V_{\mathrm{SP}}$ and the shear frequency at fixed ionic positions (classical at $T=0$ K) between the interatomic potential we used in this work44, 45, LDA, and theoretical calculations and experiments from previous works.
+The values are a little bit lower but close to that of adiabatic-connection fluctuation-dissipation theorem within the random phase approximation (ACFDT-RPA)46 and experimental data26.
 2 layers
@@
 (meV/atom)
-1.24 1.58 1.53 (RPA) Zhou et al. (2015)
+1.24 1.58 1.53 (RPA) 46
 Shear frequency
 (cm-1) (bulk)
-37 42 42-45 Lebedeva et al. (2017)
-The lower layer is fixed in position A, and there are two minima of the same energy, AB and AC, at a distance of a bond length $d=1.42$.
+37 42 42-45 26
+The lower layer is fixed in position A, and there are two minima of the same energy, AB and AC, at a distance of a bond length $d=1.42$ Å.
 For layers to slide with respect to each other, they have to overcome a barrier $V_{\mathrm{SP}}$.
+Is is worth noting that although the profile is for bilayer graphene, some of us previously showed in first-principles calculations that using 6 layers – 3 layers displaced relative to the other 3 – the profile is almost identical (see SI of Ref.
+9).
+Since interactions with subsequent layers are even weaker, the barrier should stay essentially unchanged with additional layers.
+In addition, in Ref.
+47 we obtained that energy differences between different functionals for different stacking sequences, including van-der-Waals corrections, differ by around 0.01 meV/atom.
+Thus, the barrier in the bulk limit should correspond closely to the bilayer one.
 SP is a saddle point in the full energy landscape, so the shear mode in the perpendicular direction (along the direction of two nearby AA maxima) is stable.
@@
 The lines are a guide to the eye, as in the other figures of this work.
+The standard quantum SSCHA result (blue) and its classical limit (red) give almost identical results, showing quantum effects are negligible above 100 K.
 The barrier at 100 K is reduced by a substantial 35% with respect to the value at fixed nuclei.
@@
 In the fixed ion calculation, the system is exactly at SP, but in reality, the system is also exploring surrounding configurations, which have a lower energy.
+In an analogous fashion, there is a barrier in real space, since the interlayer distance at AB is lower relative to SP.
+At 100 K, this barrier gets similarly reduced by 38%, from $0.029\,\text{\AA }$ to $0.018\,\text{\AA }$, and also shows a temperature dependence similar to that of the free energy barrier (see Figs.
+S5 and S6).
 Remarkably, our results show improved agreement with experiments.
-In a previous workNery et al.
-(2020), using density functional theory with an LDA functional (see Table 1), some of us obtained that the maximum shear stress in simulated stick-slip motion is 0.2 GPa, and we mentioned a good agreement with experimental values 0.14 GPa (Ref.
-Liu et al., 2012) –a benchmark for defect free single crystal graphiteBlomquist et al.
-(2019)– and 0.1 GPa (Ref.
-Dienwiebel et al., 2004).
-Assuming that the maximum shear stress decreases by 35% in line with the barrier, we now obtain 0.13 GPa, in excellent agreement with experimentsDienwiebel et al.
-(2004); Liu et al.
-(2012).
-[…]
+In a previous work9, using density functional theory with an LDA functional (see Table 1), some of us obtained that the maximum shear stress in simulated stick-slip motion is 0.2 GPa, and we mentioned a good agreement with experimental values 0.14 GPa (Ref.
+15) –a benchmark for defect free single crystal graphite48– and 0.1 GPa (Ref.
+23).
+Assuming that the maximum shear stress decreases by 35% in line with the barrier, we now obtain 0.13 GPa, in excellent agreement with experiments23, 15.
+Since 100 K – about 70 cm-1 in wavenumber units – is a relatively low temperature, quantum fluctuations may play a role.
+To assess their impact on the free energy barrier, we also performed calculations with the classical distribution, corresponding to the classical limit $\hbar\omega/kT\rightarrow 0$.
+In this way, $2n_{s}+1\rightarrow 2T/\omega$, and the quantum (Bose-Einstein) probability distribution for ionic configurations reduces to the classical Boltzmann distribution (an explicit expression for the distribution is included in the SI).
+We can see the SSCHA results in the classical limit in red in Fig 2, which are virtually identical to the SSCHA quantum calculations, showing that quantum effects are negligible above $100\text{\,}\mathrm{K}$.
+The reason is that shear modes responsible for the barrier reduction have very low energies, and even a modest temperature can populate such modes.
+This is not a general feature, as other systems present significant differences between classical and quantum free energy barriers, as in materials formed by light hydrogen atoms.
+Such quantum effects have been observed at 100 K (Ref.
+19), 250 K (Ref.
+21), and even higher temperatures20.
 The size of the graphene flakes considered in our work in principle correspond to nanometers up to possibly micrometer scales.
-In clean samples with micrometer sized flakes, different regions can have different stackings (AB and AC)Henck et al.
-(2018); Jiang et al.
-(2018), and a domain wall in the metastable state separates both regions.
+In clean samples with micrometer sized flakes, different regions can have different stackings (AB and AC)12, 49, and a domain wall in the metastable state separates both regions.
 The displacement of layers in this case is more complex and is not considered in our work.
 We expect shear stress to be lower in this case, since the metastable state is already present.
-Defects and stacking faults also lead to lower friction due to incommensurabilities, so the shear stress of 0.14 GPa corresponds to the maximum expected valueBlomquist et al.
-(2019).
+Defects and stacking faults also lead to lower friction due to incommensurabilities, so the shear stress of 0.14 GPa corresponds to the maximum expected value 48.
+When it comes to displacing many layers in multilayer graphene or graphite, the change of the barrier could in principle differ, since it depends on the phonon frequencies (and interlayer modes depend noticeably on the number of layers).
+For example, graphite exhibits a larger shear frequency at equilibrium, since a layer experiences primarily the restoring force of two layers instead of one (see the next subsection for more details).
+However, layers are typically not simultaneously in unstable positions relative to both nearest layers, so, as noted earlier, the barrier remains approximately unchanged in bulk.
+Since the frequency at AB increases while remaining similar at SP, the barrier reduction in bulk might be even greater than in the bilayer case.
+Other variations may arise from the choice of the interlayer interatomic potential: as shown in Table I, the RPA barrier is higher, indicating that the curvature around SP could be larger.
+Thus, vibrations around SP with more accurate interlayer potentials might yield a larger absolute barrier reduction, although the relative change may still be similar.
 In other materials, the anharmonicities associated with the unstable modes at a saddle point should also contribute negatively to the free energy.
@@
 The temperature dependence of the shear mode is similar to that of the barrier, varying by about 20% from 0 to 1000 K, and having a larger dependence at temperatures above 500 K.
-Figure 3: Bilayer graphene phonon dispersion.
-The inset shows a zoom in close to $\Gamma$ to better visualize the doubly degenerate shear modes S at $\Gamma$, and the breathing mode B.
-Due to the high speed velocity in graphene, the region of the shear mode in the BZ is small (a fraction of $\Gamma$M), and is overrepresented in small supercell calculations.
-The breathing mode, on the other hand, is flat, so nearby sampling points are not necessary to converge the free energy.
+[…]
 Figure 4: Shear frequency as a function of temperature using SSCHA (blue circles) and QHA (red crosses).
 There is no experimental data for the temperature dependence of the shear frequency as a function of temperature for bilayer graphene.
-Experimental values correspond to bulk (blue crosses) and folded 2+2 graphene (orange circles), which are corrected by the factor of a nearest neighbor modelHe et al.
-(2016) to compare to the bilayer case (and rigidly shifted by -5 cm-1 to facilitate the visual comparison).
+Experimental values correspond to bulk (blue crosses) and folded 2+2 graphene (orange circles), which are corrected by the factor of a nearest neighbor model50 to compare to the bilayer case (and rigidly shifted by -5 cm-1 to facilitate the visual comparison).
 The agreement with the temperature dependence of the bulk values is good.
 Regarding the experimental data, in Ref.
-Cong and Yu, 2014, the measurements are done in 2+2 folded graphene (4 layers).
-In order to compare with our 2-layer calculations, we multiply by a correction factor using a nearest-neighbor model for layersHe et al.
-(2016).
-In the case of bulk, such a factor is $\sqrt{2}$ (the force exerted by a layer below and above, proportional to the square of the frequency, is doubled when considering only nearest neighbor interactions), which we use to compare to the bulk measurements of Ref.
-Mishina et al., 2000.
+33 the measurements are done in 2+2 folded graphene (4 layers), and in Ref.
+34 they are done in bulk.
+To compare with our bilayer calculations, we rely on a nearest layer model which has simple conversion factors between different number of layers29.
+Since interlayer interactions in graphite are known to be weak, this should provide a good approximation.
+Indeed, there are multiple references that have measured the shear frequency for different number of layers and obtained that the model describes the trend very well, like in multilayer graphene 29 and NbSe2 50.
+The bulk case can be readily understood: while in the bilayer case the restitutive lateral force is exerted only by the other layer, in the bulk case each layer is surrounded by two of them, so the force is doubled.
+Since the interatomic force constant matrix is proportional to the square of the phonon frequencies, the shear frequency in bilayer is $\sqrt{2}$ smaller than the bulk frequency in this model.
+Using these factors29, we obtain the adjusted experimental values of Fig.
+4.
 The temperature dependence of both SSCHA and QHA curves is similar to the experimental adjusted bulk values.
@@
 The shear frequency has also been measured in other materials like bilayer NbSe2 in Ref.
-He et al., 2016, or bulk $h$-BN in Ref.
-Cuscó et al., 2019, which show a similar temperature dependence (adjusting the bulk value by a $\sqrt{2}$ factor) of roughly -0.5 cm-1/100 K.
+50, or bulk $h$-BN in Ref.
+51, which show a similar temperature dependence (adjusting the bulk value by a $\sqrt{2}$ factor) of roughly -0.5 cm-1/100 K.
 III.3 Thermal expansion
 Figure 5: (a) Graphite out-of-plane coefficient of TE as a function of temperature using SSCHA and QHA.
-PIMD calculationsHerrero and Ramírez (2020) and experimental data Bailey and Yates (1970) are also included.
+PIMD calculations39 and experimental data 37 are also included.
 PIMD overall agrees well, while QHA and SSCHA (with the potential of Ref.
-Wen et al., 2018) underestimates the TE.
+45) underestimates the TE.
 (b) Graphite in-plane coefficient of TE using SSCHA and QHA.
@@
 SSCHA gives excellent results both at low and high temperatures, considering also the disagreement within experimental data.
-To further check the adequacy of our method (the use of SSCHA with the in-planeLibbi et al.
-(2020) and out-of-planeWen et al.
-(2018) interatomic potentials), we look at the TE in the previous calculations and also in bulk graphite, and compare to experiments.
+To further check the adequacy of our method (the use of SSCHA with the in-plane44 and out-of-plane45 interatomic potentials), we look at the TE in the previous calculations and also in bulk graphite, and compare to experiments.
 The TE coefficients are given by $\alpha_{\mathrm{a}}(T)=\tfrac{1}{a}\tfrac{da}{dT}$ and $\alpha_{\mathrm{c}}(T)=\tfrac{1}{c}\tfrac{dc}{dT}$, with $a$ and $c$ the in-plane and out-of-plane lattice parameters, respectively.
 In Fig.
-5, we show the TE coefficient calculated with SSCHA and QHA in bulk, compared with experimental dataBailey and Yates (1970); Marsden et al.
-(2018) and PIMD calculations Herrero and Ramírez (2020).
-SSCHA results underestimate the TE.
-PIMD results are overall very good, although they also give values somewhat lower than experimentsBailey and Yates (1970); Marsden et al.
-(2018).
-In-plane SSCHA calculations give excellent results.
+5, we show the TE coefficient calculated with SSCHA and QHA in bulk, and compare with experimental data37, 40 and PIMD calculations 39.
+SSCHA results underestimate the TE , which we attribute to the interlayer potential45.
+This potential is fitted to energies and forces that use the many-body dispersion method in conjunction with a PBE functional, as it reproduces binding energies and interlayer distances reasonably well.
+However, it yields an elastic modulus along the c-axis about 15% lower than experimental values (see more details within Ref.
+45).
+Unsurprisingly, Ref.39 obtains better results with the LCBOPII potential52 – although the TE remains somewhat below experiments37, 40 – since it uses the experimental out of plane compressibility as a fitting parameter53.
+The bilayer TE is also similarly underestimated (see SI).
+Thus, the variation of the potential around the interlayer equilibrium distance seems to be the main factor influencing the TE, as opposed to the weaker long-range interactions involving subsequent layers.
+It is also possible that anharmonicities, captured more accurately by PIMD relative to SSCHA, play a role at high temperatures.
+In any case, for the purpose of studying the barrier – the focus of our work – the problem with the LCBOPII potential is that it includes no registry effects, making the barrier essentially non-existent.
+While the out-of-plane TE we obtained is not optimal, in-plane SSCHA calculations give excellent results.
 QHA works well at low temperatures but departs from experiments at higher temperatures.
+The negative TE in graphite and other layered systems is well-known and arises from flexural (out-of-plane) modes 54.
+We briefly summarize the mechanism here.
+In flexural modes, atoms are displaced out-of-plane, reducing the average projected in-plane distance, even though the 3D bond lengths remain essentially unchanged.
+This can also be understood intuitively via the “membrane effect”: when a layer is stretched, the out-of-plane amplitude decreases and the frequency increases, which corresponds to a negative Grüneisen parameter.
+See, for example, Ref.
+54 for more details.
 We conducted a careful convergence of these results, reported in the SI, Sec.
@@
 PIMD calculations of Ref.
-Herrero and Ramírez, 2020, although they work very well out-of-plane, greatly overestimate in-plane TE, presumably due to the inadequacy of their interatomic potential.
+39, although they work very well out-of-plane, greatly overestimate in-plane TE, presumably due to the inadequacy of their interatomic potential.
 It would be interesting in future work to compare SSCHA and PIMD using the same potentials.
@@
 We also obtained that the temperature dependence of the barrier is similar to that of the shear frequency, which agrees well with existing measurements.
+Data availability
+Additional data supporting this article has been included as part of the Supplementary Information. The SSCHA42 code used in this article is publicly available at https://sscha.eu/.
+Conflicts of interest
+There are no conflicts to declare.
 Acknowledgments
@@
 101151380.
-Supplementary Information
-S1 Barrier
-To determine the barrier, the free energy has to be determined at the equilibrium position and SP.
-At the equilibrium position, there are two shear modes, indicated with $\mathbf{S}$ in Fig.
-3(b).
-The mode perpendicular to the bond direction shown in Fig.
-1 is stable in SP, while the shear mode along the reaction coordinate (using the language of chemical reactions) is unstable.
-As opposed to the breathing mode, that is flat around $\Gamma$, the shear modes are only present in a small region of the BZ ($1/N^{2}$, with $N$ larger than 20), as a result of the large speed of sound in graphene.
-So in small supercells, the shear modes will be over-represented.
-This is more problematic at SP, where the SSCHA frequency corresponding to she unstable mode takes very small values (see Fig.
-S1), which makes the convergence of the free energy hard, mostly due to the logarithm term in $F_{\mathcal{H}}$ in Eq.
-(4).
-Just determining the barrier with this equation and a $N\times N\times 1$ supercell, results in Fig.
-S2, which is not well converged.
-All SSCHA calculations used 24000 configurations, both for the barrier and the TE.
-In order to get better convergence with $N$, we use the interpolation method mentioned in Sec.
-II.2, and obtain a much more stable barrier as a function of $1/N$ (see Fig.
-S3).
-The $40\times 40$ interpolation grid used for $F_{\mathcal{H}}$ corresponds to a supercell of 6400 atoms.
-For 100, 200 and 300 K, the barrier seems well converged.
-For a 1000 K, the trend seems to indicate convergence within 0.02 meV/atom, so for the intermediate temperatures 400 and 500 K, the barrier is likely also converged within a similar error, resulting in Fig.
-2 of the main text.
-To better understand the reduction in the barrier, we show in Fig.
-S4 the SSCHA frequencies (around the shear modes) at equilibrium and SP at $T=100$ K.
-The SSCHA frequencies at equilibrium are similar to the standard harmonic values, as shown in the main text.
-The SSCHA shear frequencies at SP instead change significantly.
-The shear mode that looks like an acoustic mode takes a value close to 2 cm-1 at $\Gamma$ (see also Fig.
-S1).
-The rest of the SSCHA phonon dispersion presents small differences between equilibrium and SP frequencies.
-In addition, we performed calculations in the classical limit ($\hbar\omega/kT\rightarrow 0$) by increasing the carbon atomic masses by a factor of 600.
-This reduces the Debye temperature below 100 K, ensuring that all modes are thermally populated and recovering the classical Boltzmann distribution for the distorted configurations.
-As we can see in Fig.
-S5, the standard quantum SSCHA calculation and its classical limit coincide, showing quantum effects are neglegible above 100 K.
-Figure S1: Unstable shear frequency at SP as a function of $1/N$, for different temperatures. The fact that it does not stabilize shows that the shear mode is present in a small region of the BZ.
-Figure S2: Barrier as a function of temperature, for different supercells, just using the $N\times N\times 1$ grid to determine the harmonic part of the free energy Eq.
-(4).
-The barrier changes significantly with $N$, which means the shear mode is overrepresented.
-On the other hand, interpolating the harmonic part of the free energy, convergence is much better.
-See Fig.
-S3.
-Figure S3: Barrier as a function of $1/N$, for different temperatures, interpolating the first term of Eq.
-4.
-The barrier seems well converged at low temperatures and harder to converge at higher temperatures.
-Values are likely converged within less than 0.05 meV/atom.
-Figure S4: SSCHA frequencies around the shear modes, at the equilibrium position (blue) and SP (red).
-Figure S5: Comparison between the standard quantum SSCHA results and its classical limit.
-S2 Shear and breathing frequency
-The SSCHA auxiliary and physical frequencies as a function of temperature can be observed in Fig.
-S6.
-The frequencies are not fully converged, and the converged value might be similar to the QHA.
-The bulk temperature dependence (corrected with the factor of the nearest neighbor model) agrees very well with QHA.
-The other mode at $\Gamma$ that involves the layers moving out of phase is the breathing mode (out-of-plane instead of in-plane).
-Its temperature dependence can be seen in Fig.
-S7.
-It decreases more at low temperatures compared to the shear mode and barrier.
-There are no measurements of the breathing mode as a function of temperature, so our calculations serve as a prediction.
-The SSCHA curve is very well converged, so we expect it to have a temperature dependence that is closer to the experiment relative to the QHA curve.
-The auxiliary frequency for $N=8$ and 10 is almost the same (Fig.
-S8), so the physical values are also converged (within 0.5 cm-1 at 1000 K, and less for lower temperatures).
-Figure S6: Shear frequency as a function of temperature, both for the physical frequency (full lines) and the auxiliary one (dashed).
-The experimental value is shifted and included twice, to facilitate the comparison of the temperature dependence with the QHA and SSCHA curves.
-The agreement of the experimental bulk value (adjusted with the nearest neighbor model to the bilayer case) with QHA is excellent, and with SSCHA is very good as well.
-Figure S7: Breathing frequency as a function of temperature using SSCHA (black circles) and QHA (red crosses).
-The SSCHA calculations are well converged and serve as a prediction, since there is no experimental data in the literature.
-Figure S8: Convergence of the auxiliary (dashed) and physical (full) breathing frequency. The physical frequency is well-converged at $N=10$.
-S3 Thermal expansion (TE)
-For the SSCHA out-of-plane TE calculations, we also used the interpolation scheme with a $40\times 40$ grid.
-The interlayer distance and in-plane lattice parameter obtained in bulk with SSCHA can be seen in Fig.
-S9.
-The fits are used to reduce numerical error, and correspond Eq.
-(6) of Ref.
-Zorzi and Perottoni, 2021,
-$b(T)=b_{0}\mathrm{exp}\left(\frac{X\theta}{e^{X/T}-1}\right),$ (S1)
-which depends on parameters $b_{0}$, $X$ and $\theta$ ($b$ can be $a$ or $c_{d}$).
-The derivative of the fits give the coefficients of TE.
-In the QHA case, the TE can be calculated quickly and accurately for many temperatures, so we just use finite differences.
-In Fig.
-S10 we compare SSCHA and QHA for bilayer graphene and bulk.
-The TE for both SSCHA and QHA bulk is lower relative to bilayer.
-In-plane, Fig.
-S11, we can also see a clear reduction of the TE coefficient from monolayer to bulk.
-In what follows, we describe in detail convergence studies we did for SSCHA, and also for QHA calculations.
-Figure S9: Out-of plane (a) and in-plane (b) SSCHA thermal expansion (blue dots). The fit corresponds to Eq. (S1).
-Figure S10: SSCHA and QHA out-of-plane coefficient of TE for graphite and bilayer graphene.
-Figure S11: In-plane coefficient of TE using SSCHA and QHA, for monolayer, bilayer and bulk.
-S3.1 Bilayer
-QHA.
-We first considered the out-of-plane TE by fixing $a$, and then we considered the full minimization of the free energy by varying both $a$ and $c_{d}$.
-In Fig.
-S12 we can see $c_{d}(T)$ both for fixed $a$, and for the full minimization.
-The full minimization has a slightly larger dependence with temperature, but fixing $a$ is a good approximation.
-Fig.
-S13 is analogous for $a(T)$, with a fixed value $c_{d}=3.5$.
-The curves are very similar, except for a small difference at high temperatures (of about 0.0005 Å at 1000 K).
-We also studied convergence with the size of the grid (Fig.
-S14).
-At $N=12$, $c_{d}(T)$ is well converged (within 0.005 Å at 1000 K), although $N\sim 20$ to achieve full convergence.
-The shear mode increases TE because when layers are displaced relative to each other, the interlayer distance is larger (see Fig.
-S15).
-The monotonous behavior indicates that the shear modes are overrepresented for coarser grids, and become better sampled with denser grids (having wavevectors $\mathbf{q}$ closer to $\Gamma$).
-In-plane, a larger $N$ of about 50 is needed to fully achieve convergence (Fig.
-S16).
-Again, monotonous convergence suggests inadequate sampling close to $\Gamma$.
-Since the convergence in bulk is much faster (see below Fig.
-S24), we attribute the slow convergence to the flexural mode, which becomes linear (an acoustic mode) and easier to sample in bulk.
-[…]
-On the other hand, $a(T)$ has less variation with temperature relative to QHA and is already converged at about $N=8$ (Fig. S18).
-S3.2 Bulk
-QHA.
-In bulk, there is virtually no difference for the out-of-plane TE when fixing $a$, or considering also the minimization of $a$ (Fig.
-S20).
-In-plane (Fig.
-S21), there is a small difference between both fixing or varying $a$, similar to the bilayer case.
-Just like in the bilayer case, relative dense grids are needed to achieve out-of-plane convergence (Fig.
-S22), which is again likely due to the shear mode.
-Regarding the extension of the supercell in the out-of-plane direction, 2 unit cells (4 layers) give good results (Fig.
-S23).
-In-plane (Fig.
-S24), convergence is much faster than out-of-plane.
-SSCHA. Here we also used the interpolation method ($40\times 40$ grid), to go beyond the $N=8$ supercell that was possible in the direct SSCHA calculation, resulting in the blue dots of Fig. S9 (a).
-In-plane, based on $N=8$ already giving converged results in bilayer (Fig.
-S18), and also considering that the QHA convergence in bulk (Fig.
-S24) is much faster than in bilayer (Fig.
-S16), $N=8$ should give converged SSCHA results in bulk, which corresponds to the blue dots in Fig.
-S9 (b).
-Figure S12: Interlayer distance $c_{d}$ as a functi on of temperature in bilayer graphene using QHA, fixing the value of $a$ (dashed-blue) and doing the full minimization (full-black).
-In-plane TE has little effect on the out-of-plane TE.
-Figure S13: Lattice parameter $a$ as a function of temperature in bilayer graphene using QHA, fixing the value of $c_{d}$ (dashed-blue) and doing the full minimization (full-black).
-In-plane TE has little effect on the out-of-plane TE.
-Figure S14: Interlayer distance as a function of temperture in bilayer graphene, using QHA. At $N\sim 20$ $c_{d}$ is converged, although $N=12$ also gives a good result.
-Figure S15: Interlayer distance profile along the bond direction, when moving a graphene layer relative to another one (analogous to Fig.
-1, but for distance instead of energy).
-The minimum distance is at the equilibrium positions.
-Figure S16: In-plane lattice parameter as a function of temperature in bilayer graphene for different $N$ using QHA. A large $N$ is needed to achieve convergence.
-Figure S17: Interlayer distance as a function of temperature in bilayer graphene for different $N$ using SSCHA.
-Convergence is similar to the QHA case, Fig.
-S14, where $N=12$ gives a good value of $c_{d}(T)$.
-Figure S18: In-plane lattice parameter as a function of temperature, for different $N$. At $N=8$ the curve is already well converged.
-Figure S19: Interlayer distance as a function of temperature in bilayer graphene, using SSCHA, at two different values of the in-plane lattice parameter.
-The curves are similar to the ones of Fig.
-S12.
-In-plane TE has little effect on the out-of-plane TE.
-Figure S20: Interlayer distance as a function of temperature in bulk using QHA, fixing the value of $c_{d}$ (dashed-blue) and doing the full minimization (full-black).
-In-plane TE has barely no effect on the out-of-plane TE.
-Figure S21: In-plane lattice parameter as a function of temperature in bulk using QHA, fixing the value of $c_{d}$ (dashed-blue) and doing the full minimization (full-black).
-The out-of-plane TE has a slightly larger but similar effect relative to bilayer on the in-plane TE.
-Figure S22: Interlayer distance as a function of temperature in graphite for different $N$ using QHA. Convergence is similar to the bilayer case.
-Figure S23: Interlayer distance as a function of temperature in graphite for different $N_{z}$. $N_{z}=2$ is the value used in our calculations.
-Figure S24: In-plane lattice parameter as a function of temperature in graphite using QHA, for different $N$. At $N=6$ convergence is already very good, as opposed to the bilayer case Fig. S17.